### v1

# Jarník-Besicovitch type theorems for semisimple algebraic groups

Abstract.
In this note, we initiate a study on Jarník-Besicovitch type theorems for semisimple algebraic groups from the representation-theoretic point of view. Let $\rho:\mathbf{G}\to\operatorname{\mathbf{GL}}(V)$ be an irreducible $\mathbb{Q}$-rational representation of a connected semisimple $\mathbb{Q}$-algebraic group $\mathbf{G}$ on a complex vector space $V$ and $\{a_{t}\}_{t\in\mathbb{R}}$ a one-parameter subgroup in a $\mathbb{Q}$-split torus in $\mathbf{G}$. We define a subset $S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})$ of Diophantine elements of type $\tau$ in $\mathbf{G}(\mathbb{R})$ in terms of the representation $\rho$ and the subgroup $\{a_{t}\}_{t\in\mathbb{R}}$, and prove formulas for the Hausdorff dimension of the complement of $S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})$. As corollaries, we deduce several Jarník-Besicovitch type theorems.
2020 Mathematics Subject Classification:
Primary: 37A17; Secondary: 11J83
The author acknowledges the support by the Institute of Modern Analysis-A Frontier Research Centre of Shanghai.
1. Introduction and main results
1.1. Introduction
In this note, we initiate a study on the Jarník-Besicovitch theorem in the metric theory of Diophantine approximation. Recall that a real number $x\in\mathbb{R}$ is Diophantine of type $\kappa>0$ if there exists a constant $C>0$ such that
$\left|nx-m\right|\geq\frac{C}{n^{\kappa}}\;\left(\forall\frac{m}{n}\in\mathbb{Q}\right).$
Then necessarily $x\in\mathbb{R}\setminus\mathbb{Q}$ and the Dirichlet theorem implies that $\kappa\geq 1$. Denote by $S_{\kappa}$ the set of all Diophantine numbers of type $\kappa$ and $S_{\kappa}^{c}$ its complement. Then one can deduce from the Khintchine’s theorem on metric Diophantine approximation that $S_{\kappa}^{c}$ is a set of full measure if $\kappa=1$, and is null if $\kappa>1$. A refined result due to Jarník and Besicovitch [4, 22] states further that the Hausdorff dimension of $S_{\kappa}^{c}$ is equal to $2/(\kappa+1)$ ($\kappa>1$).
From the viewpoint of dynamical systems, the Jarník-Besicovitch theorem can be reformulated as a result of shrinking target problem, which was first observed by Hill and Velani [20]. Let $f:X\to X$ be a map on a metric space $X$ with a measure $\mu$. Generally speaking, in the shrinking target problem, one studies the set $S$ of points in $X$ whose trajectories under $f$ approach to a shrinking target infinitely often with certain rate, and seeks to establish results about the size (the measure or the Hausdorff dimension) of the set $S$. The measure theoretic version of the shrinking target problem is usually related to the Khintchine’s theorem [26, 41] while the Hausdorff dimension version is linked with the Jarník-Besicovitch theorem [20].
In homogeneous dynamics, various Diophantine sets can also be studied by the dynamical and ergodic properties of group actions on homogeneous spaces via Dani’s correspondence [14, 26]. Specifically, let $X_{m+n}=\operatorname{\mathbf{SL}}_{m+n}(\mathbb{R})/\operatorname{\mathbf{SL}}_{m+n}(\mathbb{Z})$,
$a_{t}=\operatorname{diag}(e^{t/m},\dots,e^{t/m},e^{-t/n},\dots,e^{-t/n})\quad(t\in\mathbb{R})$
[…]
$g\operatorname{\mathbf{SL}}_{m+n}(\mathbb{Z})\to g\cdot\mathbb{Z}^{m+n}\quad(g\in\operatorname{\mathbf{SL}}_{m+n}(\mathbb{R})).$
Let $\delta:X_{m+n}\to\mathbb{R}_{+}$ be the systole function on $X_{m+n}$ defined by
$\delta(\Lambda)=\inf_{v\in\Lambda\setminus\{0\}}\|v\|$
which is also called the first minimum of the lattice $\Lambda\in X_{m+n}$. For any $m\times n$ matrix $A$, denote by
$u_{A}:=\left(\begin{array}[]{ccc}I_{m}&A\\ O&I_{n}\end{array}\right)\in N_{+}.$
Then it is well-known via Mahler’s compactness criterion and Dani’s correspondence [14, 26] that $A$ is badly approximable if and only if the orbit $\{a_{t}u_{A}\cdot\mathbb{Z}^{m+n}:t\geq 0\}$ is bounded in $X_{m+n}$, i.e.
$\delta(a_{t}u_{A}\cdot\mathbb{Z}^{m+n})\geq c\quad(t\geq 0)$
for some constant $c>0$; singular if and only if $\{a_{t}u_{A}\cdot\mathbb{Z}^{m+n}:t\geq 0\}$ diverges i.e.
$\delta(a_{t}u_{A}\cdot\mathbb{Z}^{m+n})\to 0\quad(\textup{as }t\to\infty);$
and very well approximable if and only if
$\limsup_{t\to\infty}\frac{-\log(\delta(a_{t}u_{A}\cdot\mathbb{Z}^{m+n}))}{t}>0.$
We say that an $m\times n$ matrix $A$ is Diophantine of type $\kappa$ if there exists a constant $C>0$ such that
$\|Aq-p\|\geq C\|q\|^{-\kappa}$
for any $q\in\mathbb{Z}^{n}\setminus\{0\}$ and $p\in\mathbb{Z}^{m}$. Then necessarily $\kappa\geq n/m$, and again using Dani’s correspondence, one can deduce that $A$ is Diophantine of type $\kappa$ if and only if there exists a constant $C>0$ such that
$\delta(a_{t}\cdot(u_{A}\cdot\mathbb{Z}^{m+n}))\geq Ce^{-\tau t}\;(\forall t>0)$
where
$\tau=\frac{m\kappa-n}{mn(\kappa+1)}.$
The set $S_{\kappa}$ of all Diophantine numbers of type $\kappa$ defined at the beginning of the introduction then corresponds to the case $m=n=1$ here. In particular, one can rephrase the Jarník-Besicovitch theorem for the homogeneous space $X_{m+n}=X_{2}$ $(m=n=1)$, and get that for any parameter $10\leq\tau<$ 1, the Hausdorff dimension of the complement of the set
$\{u_{A}\in N_{+}:\textup{there exists $C>0$ such that }\delta(a_{t}\cdot(u_{A}\cdot\mathbb{Z}^{2}))\geq Ce^{-\tau t}\;(\forall t>0)\}$
is equal to $1-\tau$. Using the decomposition of $X_{2}$ into the stable, central and unstable submanifolds of the flow $\{a_{t}\}_{t\in\mathbb{R}}$ on $X_{2}$, one can then obtain the following equivalent statement: the Hausdorff dimension of the complement of the set
$\{p\in X_{2}:\textup{there exists $C>0$ such that }\delta(a_{t}\cdot p)\geq Ce^{-\tau t}\;(\forall t>0)\}$
is equal to $3-\tau$. In general, the matrix version of the Jarnik-Besicovitch theorem has also been established [9, 15], and it is equivalent to the statement that for any $0\leq\tau<1/n$, the Hausdorff dimension of the complement of the set
$\{p\in X_{m+n}:\textup{there exists $C>0$ such that }\delta(a_{t}\cdot p)\geq Ce^{-\tau t}\;(\forall t>0)\}$
is equal to $mn(1-\tau)+m^{2}+n^{2}+mn-1$. We remark that one can further consider the case where $\{a_{t}\}_{t\in\mathbb{R}}$ is a generic one-parameter diagonal subgroup in $\operatorname{\mathbf{SL}}_{m+n}(\mathbb{R})$ and study the Hausdorff dimension of the complement of the set
$\{p\in X_{m+n}:\textup{there exists $C>0$ such that }\delta(a_{t}\cdot p)\geq Ce^{-\tau t}\;(\forall t>0)\}.$
This type of question leads to a general weighted multidimensional Jarník-Besicovitch theorem on metric Diophantine approximation, which is one of the main topics we pursue in this note (See Theorem 1.6).
[…]
$\delta(\Lambda)=\inf_{v\in\Lambda\setminus\{0\}}\|v\|.$
Let $\{a_{t}\}_{t\in\mathbb{R}}$ be a one-parameter Ad-diagonalizable subgroup in $\mathbf{G}(\mathbb{R})$. Then the main object we would like to study is the Hausdorff dimension of the complement of the set
$\{g\in\mathbf{G}(\mathbb{R}):\delta(\rho(a_{t}\cdot g)\cdot\mathbb{Z}^{d})\geq Ce^{-\tau t}\textup{ for some }C>0\}.$
Clearly the discussions above about the Jarnik-Besicovitch theorem with the homogeneous flow $a_{t}=\operatorname{diag}(e^{t/m},\dots,e^{t/m},e^{-t/n},\dots,e^{-t/n})$ on $X_{m+n}$ can be reduced to this setting if we let $\mathbf{G}=\operatorname{\mathbf{SL}}_{m+n}$ and $\rho$ the standard representation of $\operatorname{\mathbf{SL}}_{m+n}$ on $\mathbb{C}^{m+n}$. Moreover, we will see later (Theorem 1.8) that by considering representations of algebraic groups, one will be able to establish some Jarnik-Besicovitch type theorems for certain algebraic subvarieties in an affine space. We remark that the question we study in this note starts from the sparse equidistribution problem, and a related shrinking target problem plays an important role. For more details, one can read [17, 46, 47, 48].
1.2. Main results
In the rest of this note, we address the question above under the following assumption. We assume that $\mathbf{G}$ is a connected semisimple $\mathbb{Q}$-algebraic group (in the Zariski topology) and $\mathbf{T}$ is a maximal $\mathbb{Q}$-split torus in $\mathbf{G}$. Let $\{a_{t}\}_{t\in\mathbb{R}}$ be a one-parameter subgroup in $\mathbf{T}(\mathbb{R})$, and $\rho$ a finite-dimensional irreducible representation of $\mathbf{G}$ defined over $\mathbb{Q}$ on a complex vector space $V$ with $\dim\ker\rho=0$. Let $d=\dim V$ and we may identify $V\cong\mathbb{C}^{d}$ so that $\mathbb{Z}^{d}\subset\mathbb{C}^{d}$ is compatible with the $\mathbb{Q}$-structure in $V$.
Definition 1.1.
An element $g\in\mathbf{G}(\mathbb{R})$ is Diophantine of type $\tau\geq 0$ if there exists a constant $C>0$ such that
$\delta(\rho(a_{t}\cdot g)\mathbb{Z}^{d})\geq Ce^{-\tau t}\textup{ for any }t>0$
where $\delta$ is the first minimum function. We denote by $S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})$ the set of Diophantine elements of type $\tau$ in $\mathbf{G}(\mathbb{R})$ and $S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}$ its complement in $\mathbf{G}(\mathbb{R})$. If the representation $\rho$ and the one-parameter subgroup $\{a_{t}\}_{t\in\mathbb{R}}$ are clearly stated in the contexts, we will simply write $S_{\tau}$ instead of $S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})$.
Remark 1.2.
If $\mathbf{G}$ is $\mathbb{Q}$-anisotropic, then $\mathbf{T}=\{e\}$. So in the following, we assume that $\mathbf{G}$ is $\mathbb{Q}$-isotropic.
To state the first main theorem about $\dim_{H}S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}$, we need to introduce some notation. We choose a minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$ in $\mathbf{G}$ containing $\mathbf{T}$. Then $\mathbf{P}_{0}$ and $\mathbf{T}$ defines a root system ($\Phi,\;\Phi^{+},\;\Delta$) where $\Phi$ is the set of $\mathbb{Q}$-roots relative to $\mathbf{T}$, $\Phi^{+}$ is the set of positive $\mathbb{Q}$-roots determined by $\mathbf{P}_{0}$ and $\Delta$ is the set of simple $\mathbb{Q}$-roots in $\Phi^{+}$. Let $\overline{\mathbf{P}}_{0}$ be the opposite minimal parabolic $\mathbb{Q}$-subgroup of $\mathbf{P}_{0}$ defined by $\Phi\setminus\Phi^{+}$. Without loss of generality, we may assume that the stable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in the unipotent radical $R_{u}(\mathbf{P}_{0})$ of $\mathbf{P}_{0}$ and its unstable horospherical subgroup is contained in the unipotent radical $R_{u}(\overline{\mathbf{P}}_{0})$ of $\overline{\mathbf{P}}_{0}$. One can write the space $V$ in the representation $\rho$ as a direct sum of weight spaces with respect to the action of $\mathbf{T}$
$V=\bigoplus_{\beta}V_{\beta}.$
By the structure theory of irreducible representations of complex semisimple groups and semisimple Lie algebras, there is a highest weight $\beta_{0}$ among the weights $\beta$’s (where the order is determined by the minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$) and we denote by $V_{\beta_{0}}$ its corresponding weight space. (See §2 for more details). The stabilizer of the weight space $V_{\beta_{0}}$ in $\mathbf{G}$ is a parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{\beta_{0}}$ containing $\mathbf{P}_{0}$, and its unipotent radical is denoted by $R_{u}(\mathbf{P}_{\beta_{0}})$. The opposite parabolic $\mathbb{Q}$-subgroup of $\mathbf{P}_{\beta_{0}}$ containing $\overline{\mathbf{P}}_{0}$ and its unipotent radical are denoted by $\overline{\mathbf{P}}_{\beta_{0}}$ and $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ respectively.
In the following, if an algebraic $\mathbb{Q}$-subgroup $\mathbf{F}\subset\mathbf{G}$ is normalized by $\mathbf{T}$, then we write $\Phi(\mathbf{F})$ for the set of $\mathbb{Q}$-roots in $\mathbf{F}$ relative to $\mathbf{T}$ while the summation $\sum_{\alpha\in\Phi(\mathbf{F})}$ (or $\prod_{\alpha\in\Phi(\mathbf{F})}$) means that we take the sum (or product) all over $\alpha\in\Phi(\mathbf{F})$ counted with multiplicities (i.e., the dimensions of the corresponding $\mathbb{Q}$-root spaces in the Lie algebra of $\mathbf{F}$). For a $\mathbb{Q}$-root or a $\mathbb{Q}$-weight $\lambda$ in $\rho$ relative to $\mathbf{T}$ (which is a $\mathbb{Q}$-character of $\mathbf{T}$) , we will often consider it as a linear functional on the Lie algebra $\mathfrak{a}$ of $\mathbf{T}(\mathbb{R})$ and use the same symbol. In particular, we will write $\lambda(a_{t})$ ($t\in\mathbb{R}$) for the values of $\lambda$ (as a linear functional) on the Lie algebra of $\{a_{t}\}_{t\in\mathbb{R}}$ (so that $\lambda(a_{t})$ is linear on $t\in\mathbb{R}$). We denote by $\nu_{0}$ the $\mathbb{Q}$-root in $R_{u}(\overline{\mathbf{P}}_{0})$ relative to $\mathbf{T}$ such that
$\nu_{0}(a_{1})=\max\{\alpha(a_{1}):\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{0}))\}.$
Now we can state the first main theorem in this note.
Theorem 1.3.
Let $\mathbf{G}$ be a connected semisimple algebraic group defined over $\mathbb{Q}$, $\mathbf{T}$ a maximal $\mathbb{Q}$-split torus in $\mathbf{G}$ and $\{a_{t}\}_{t\in\mathbb{R}}$ a one-parameter subgroup in $\mathbf{T}(\mathbb{R})$. Let $\rho$ be a finite-dimensional irreducible representation of $\mathbf{G}$ defined over $\mathbb{Q}$ on a complex vector space $V$ with $\dim\ker\rho=0$. Then
$\dim_{H}S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}\geq$ $\dim\mathbf{G}-\frac{\tau}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}\alpha(a_{1})$
for any $0\leq\tau<\beta_{0}(a_{-1})$.
To state the next theorem, we need to introduce another notation. Let $N(\mathbf{T})$ and $Z(\mathbf{T})$ be the normalizer and centralizer of $\mathbf{T}$ respectively. Then the Weyl group relative to $\mathbb{Q}$ is defined by
${}_{\mathbb{Q}}{W}=N(\mathbf{T})/Z(\mathbf{T}).$
Let $\mathbf{P}_{0}$ be the minimal parabolic $\mathbb{Q}$-subgroup of $\mathbf{G}$ as defined above, and we may assume that the stable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in the unipotent radical $R_{u}(\mathbf{P}_{0})$ of $\mathbf{P}_{0}$ and its unstable horospherical subgroup is contained in the unipotent radical $R_{u}(\overline{\mathbf{P}}_{0})$ of $\overline{\mathbf{P}}_{0}$. The Bruhat decomposition of $\mathbf{G}$ is the following [6, §21]
$\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot\mathbf{P}_{0}(\mathbb{Q})$
which implies that
$\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot\overline{\mathbf{P}}_{0}(\mathbb{Q}),\quad\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{Q}).$
Let $\{w_{i}\}_{i\in I}$ be a set of representatives of ${}_{\mathbb{Q}}{W}$ in $\mathbf{G}(\mathbb{Q})$ where $I$ is the index set of ${}_{\mathbb{Q}}W$. For each $w_{i}$ $(i\in I)$, define
$\mathbf{F}_{w_{i}}=R_{u}(\overline{\mathbf{P}}_{0})\cap w_{i}^{-1}R_{u}(\overline{\mathbf{P}}_{\beta_{0}})w_{i},\quad\mathbf{H}_{w_{i}}=R_{u}(\overline{\mathbf{P}}_{0})\cap w_{i}^{-1}\mathbf{P}_{\beta_{0}}w_{i}$
and we have
$R_{u}(\overline{\mathbf{P}}_{0})=\mathbf{H}_{w_{i}}\cdot\mathbf{F}_{w_{i}}\quad(i\in I).$
Then one can further write the Bruhat decomposition as
$\mathbf{G}(\mathbb{Q})=$ $\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{Q})$
$=$ $\bigcup_{i\in I}\mathbf{P}_{0}(\mathbb{Q})\cdot(w_{i}\mathbf{H}_{w_{i}}w_{i}^{-1})(\mathbb{Q})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{Q})$
$=$ $\bigcup_{i\in I}\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{Q}).$
Note that the sets $\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{Q})$ $(i\in I)$ in the decomposition of $\mathbf{G}(\mathbb{Q})$ above may overlap, and for our purpose, we can choose any subset $\bar{I}$ of $I$ (as small as possible) such that
$\mathbf{G}(\mathbb{Q})=\bigcup_{i\in\bar{I}}\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{Q}).$
In the statement of Theorem 1.4 below, we will fix any such subset $\bar{I}$ in $I$.
Let $\bigwedge^{\dim V_{\beta_{0}}}V$ be the $\dim V_{\beta_{0}}$-exterior product vector space of $V$ over $\mathbb{C}$, and $\rho_{\beta_{0}}$ the natural extension of $\rho$ on $\bigwedge^{\dim V_{\beta_{0}}}V$. Let
$\{e_{1},e_{2},\dots,e_{\dim V_{\beta_{0}}}\}\subset\mathbb{Z}^{d}$
be an integral basis in $V_{\beta_{0}}$ which spans $V_{\beta_{0}}\cap\mathbb{Z}^{d}$. We write
$e_{V_{\beta_{0}}}:=e_{1}\wedge e_{2}\wedge\cdots\wedge e_{\dim V_{\beta_{0}}}\in\bigwedge^{\dim V_{\beta_{0}}}V.$
For each $i\in\bar{I}$, define the following morphism
$\Psi_{w_{i}}:w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}(\mathbb{R})\to\bigwedge^{\dim V_{\beta_{0}}}V,\quad\Psi_{w_{i}}(x)=\rho_{\beta_{0}}(x)\cdot e_{V_{\beta_{0}}}.$
Note that $w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}\subset R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ and $\Psi_{w_{i}}$ is an isomorphism onto its image. We denote by $a_{w_{i}}$ the growth rate of the asymptotic volume estimate of the real variety $\textup{Im}(\Psi_{w_{i}})$, and by $A_{w_{i}}$ the growth rate of the number of rational points in $w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}(\mathbb{Q})$ (See §6 Corollary 6.8 and equation ($\ast$6) for more details). Let $\nu_{0}$ be the $\mathbb{Q}$-root in $R_{u}(\overline{\mathbf{P}}_{0})$ such that
$\nu_{0}(a_{1})=\max\{\alpha(a_{1}):\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{0}))\}.$
Now we can state the second main theorem in this note.
Theorem 1.4.
Let $\mathbf{G}$ be a connected semisimple algebraic group defined over $\mathbb{Q}$, $\mathbf{T}$ a maximal $\mathbb{Q}$-split torus in $\mathbf{G}$ and $\{a_{t}\}_{t\in\mathbb{R}}$ a one-parameter subgroup in $\mathbf{T}(\mathbb{R})$. Let $\rho$ be a finite-dimensional irreducible representation of $\mathbf{G}$ defined over $\mathbb{Q}$ on a complex vector space $V$ with $\dim\ker\rho=0$. Then
$\dim_{H}S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}$
$\leq$ $\max_{i\in\bar{I}}\left\{\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{a_{w_{i}},A_{w_{i}}\}\cdot\dim V_{\beta_{0}}\Big{|}\beta_{0}(w_{i}a_{-1}w_{i}^{-1})>\tau\right\}$
for any $0\leq\tau<\beta_{0}(a_{-1})$.
Remark 1.5.
It follows from the proof of Theorem 1.4 that for any $\tau\geq\beta_{0}(a_{-1})$, the subset $S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}=\emptyset$. (See §6.)
As corollaries of Theorems 1.3 and 1.4, we obtain the following results. The first result generalizes the Jarnik-Besicovitch type theorem on the homogneous space $X_{n}=\operatorname{\mathbf{SL}}_{n}(\mathbb{R})/\operatorname{\mathbf{SL}}_{n}(\mathbb{Z})$ for any one-parameter diagonal flow.
Theorem 1.6.
Let $\rho:\operatorname{\mathbf{SL}}_{n}\to\operatorname{\mathbf{GL}}(V)$ be the standard representation of $\mathbf{G}=\operatorname{\mathbf{SL}}_{n}$ on the complex vector space $V=\mathbb{C}^{n}$ defined by
$\rho(g)\cdot v=g\cdot v\quad(g\in\operatorname{\mathbf{SL}}_{n},v\in V)$
via matrix multiplication. Let $\{a_{t}\}_{t\in\mathbb{R}}$ be a one-parameter diagonal subgroup in $\operatorname{\mathbf{SL}}_{n}(\mathbb{R})$, $\beta_{0}$ the highest weight of $\rho$ with respect to $\{a_{t}\}_{t\in\mathbb{R}}$ defined as in Theorem 1.3, and $\nu_{0}$ the $\mathbb{Q}$-root defined as in Theorem 1.4. Then for any $0\leq\tau<\beta_{0}(a_{-1})$, we have
$\dim_{H}S_{\tau}^{c}=\dim\mathbf{G}-\frac{n\cdot\tau}{\nu_{0}(a_{1})}.$
The second result generalizes the main results in [17].
Theorem 1.7.
Let $\rho:\mathbf{G}\to\operatorname{\mathbf{GL}}(V)$ where $\mathbf{G}=\operatorname{\mathbf{SL}}_{n}$, $V=\mathfrak{g}=\mathfrak{sl}_{n}$ is the Lie algebra of $\operatorname{\mathbf{SL}}_{n}$ and $\rho=\operatorname{Ad}$ is the adjoint representation of $\mathbf{G}$. Let $\{a_{t}\}_{t\in\mathbb{R}}$ be a one-parameter diagonal subgroup in $\mathbf{G}$, and $\nu_{0}$ the $\mathbb{Q}$-root defined as in Theorem 1.4. Then for any $0\leq\tau<\nu_{0}(a_{1})$, we have
$\dim_{H}S_{\tau}^{c}=\dim\mathbf{G}-\frac{(n-1)\cdot\tau}{\nu_{0}(a_{1})}.$
We also deduce the following
Theorem 1.8.
Let $\rho:\operatorname{\mathbf{SL}}_{2}\to\operatorname{\mathbf{GL}}(V)$ be an irreducible representation of $\operatorname{\mathbf{SL}}_{2}$ defined over $\mathbb{Q}$. Let $\{a_{t}\}_{t\in\mathbb{R}}$ be a one-parameter diagonal subgroup in $\operatorname{\mathbf{SL}}_{2}(\mathbb{R})$, $\beta_{0}$ the highest weight of $\rho$ with respect to $\{a_{t}\}_{t\in\mathbb{R}}$ defined as in Theorem 1.3, and $\nu_{0}$ the $\mathbb{Q}$-root in $\operatorname{\mathbf{SL}}_{2}$ defined as in Theorem 1.4. Then for any $0\leq\tau<\beta_{0}(a_{-1})$, we have
$\dim_{H}S_{\tau}^{c}=3-\frac{\tau}{\beta_{0}(a_{-1})}.$
Remark 1.9.
Let $\rho:\operatorname{\mathbf{SL}}_{2}\to\operatorname{\mathbf{GL}}(V)$ be the irreducible representation in Theorem 1.8 with $\dim V=n+1$ $(n\geq 1)$, $\{a_{t}\}_{t\in\mathbb{R}}$ a one-parameter diagonal subgroup in $\operatorname{\mathbf{SL}}_{2}(\mathbb{R})$ and let $\rho_{0}$ be the standard representation of $\operatorname{\mathbf{SL}}_{n+1}$ on $\mathbb{C}^{n+1}$. Note that $V\cong\mathbb{C}^{n+1}$ and $\rho(\operatorname{\mathbf{SL}}_{2})\subset\operatorname{\mathbf{SL}}_{n+1}\subset\operatorname{\mathbf{GL}}(V)$. Then by Theorem 1.8, we have
$\dim_{H}S_{\tau}(\rho_{0},\{\rho(a_{t})\}_{t\in\mathbb{R}})^{c}\cap\rho(\operatorname{\mathbf{SL}}_{2}(\mathbb{R}))=\dim_{H}S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}=3-\frac{\tau}{\beta_{0}(a_{-1})}$
where $\beta_{0}$ is the highest weight of $\rho$ with respect to $\{a_{t}\}_{t\in\mathbb{R}}$ defined as in Theorem 1.3. On the other hand, by Theorem 1.6
$\dim_{H}S_{\tau}(\rho_{0},\{\rho(a_{t})\}_{t\in\mathbb{R}})^{c}=\dim\operatorname{\mathbf{SL}}_{n+1}-\frac{(n+1)\cdot\tau}{\nu_{0}(\rho(a_{1}))}$
where $\nu_{0}$ is the $\mathbb{Q}$-root in $\operatorname{\mathbf{SL}}_{n+1}$ with respect to $\{\rho(a_{t})\}_{t\in\mathbb{R}}$ defined as in Theorem 1.4. This gives an example of Jarník-Besicovitch type theorems for some algebraic subvarieties in affine spaces.
Remark 1.10.
One can see from the arguments in this note that all the theorems stated above also work for any open bounded subset $U$ in $\mathbf{G}(\mathbb{R})$, i.e. $\dim_{H}S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}\cap U=\dim_{H}S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}$. (See Remark 5.6.)
1.3. Discussions and open problems
[…]
$\liminf_{t\to\infty}-\frac{1}{t}\log\delta(\rho(a_{t}\cdot g)\cdot\mathbb{Z}^{d})=\tau.$
When $\mathbf{G}=\operatorname{\mathbf{SL}}_{n}$ and $\rho$ is the standard representation of $\mathbf{G}$, the Hausdorff dimension and the packing dimension of $\operatorname{Sing}(\tau)$ are discussed in [43] in the case that $\{a_{t}\}_{t\in\mathbb{R}}$ is a singular one-parameter diagonal subgroup in $\operatorname{\mathbf{SL}}_{n}(\mathbb{R})$, and the proofs rely on a variational principle in the parametric geometry of numbers. This variational principle is later generalized in [39] for any one-parameter diagonal subgroup in $\operatorname{\mathbf{SL}}_{n}(\mathbb{R})$.
For the lower bound and the upper bound we obtain in Theorem 1.3 and Theorem 1.4, we have computed in several cases (e.g. Theorems 1.61.7 and 1.8) that these two bounds are equal. It would be interesting to compute these two bounds in other cases and determine the representations $\rho$ for which these two bounds are equal.
2. Preliminaries
In this section, we list some preliminaries needed in this note. Then we give a definition of rational elements in $\mathbf{G}(\mathbb{R})$ and discuss some properties of rational elements.
We first need the reduction theory of arithmetic subgroups of $\mathbf{G}(\mathbb{R})$ [5]. Let $K$ be a maximal compact subgroup in $\mathbf{G}(\mathbb{R})$ and $\Gamma$ an arithmetic subgroup in $\mathbf{G}(\mathbb{Z})$. We fix a minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$ in $\mathbf{G}$ containing $\mathbf{T}$. Denote by $\Phi$ the set of $\mathbb{Q}$-roots with respect to $\mathbf{T}$, $\Phi^{+}$ the set of positive $\mathbb{Q}$-roots corresponding to the minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$ and $\Delta$ the set of simple $\mathbb{Q}$-roots in $\Phi^{+}$. For any parabolic $\mathbb{Q}$-subgroup $\mathbf{P}$, denote by $R_{u}(\mathbf{P})$ its unipotent radical. Let $M$ be the connected component of identity in the unique maximal $\mathbb{Q}$-anisotropic subgroup in $Z_{\mathbf{G}(\mathbb{R})}(\mathbf{T}(\mathbb{R}))$ (= the centralizer of $\mathbf{T}(\mathbb{R})$ in $\mathbf{G}(\mathbb{R})$). For $\eta>0$, denote by
$T_{\eta}=\{a\in\mathbf{T}(\mathbb{R}):\lambda(a)\leq\eta,\;\lambda\textup{ a simple root in }\Delta\}.$
A Siegel set in $\mathbf{G}(\mathbb{R})$ is a subset of the form $S_{\eta,\Omega}=K\cdot T_{\eta}\cdot\Omega$ for some $\eta>0$ and a relatively compact open subset $\Omega$ containing identity in $M\cdot R_{u}(\mathbf{P}_{0})(\mathbb{R})$, and the group $\mathbf{G}(\mathbb{R})$ can be written as
$\mathbf{G}(\mathbb{R})=S_{\eta,\Omega}\cdot\mathcal{K}\cdot\Gamma$
for some Siegel set $S_{\eta,\Omega}$ and some finite subset $\mathcal{K}\subset\mathbf{G}(\mathbb{Q})$. Moreover, the finite set $\mathcal{K}$ satisfies the property that
$\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot\mathcal{K}\cdot\Gamma$
where $\mathbf{P}_{0}$ is the minimal parabolic $\mathbb{Q}$-subgroup. Denote by
$\mathcal{K}=\{x_{1},x_{2},\dots,x_{k}\}=\{x_{j}\}_{j\in J}\subset\mathbf{G}(\mathbb{Q})$
and we may assume that $e\in\mathcal{K}$. In what follows, we choose $\Gamma$ to be an arithmetic subgroup in $\mathbf{G}(\mathbb{Z})\cap\mathbf{G}(\mathbb{R})^{0}$, where $\mathbf{G}(\mathbb{R})^{0}$ denotes the connected component of identity in the Lie group $\mathbf{G}(\mathbb{R})$, so that $\rho(\Gamma)$ preserves the lattice $\mathbb{Z}^{d}$ in $V$. Without loss of generality, we may assume that the stable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in $R_{u}(\mathbf{P}_{0})$ and its unstable horospherical subgroup is contained in $R_{u}(\overline{\mathbf{P}}_{0})$ where $\overline{\mathbf{P}}_{0}$ is the opposite minimal parabolic $\mathbb{Q}$-subgroup of $\mathbf{P}_{0}$ determined by $\Phi\setminus\Phi^{+}$. We also write $\mathbf{P}_{0}=\mathbf{M}_{0}\cdot R_{u}(\mathbf{P}_{0})$ where $\mathbf{M}_{0}$ is the centralizer of $\mathbf{T}$ in $\mathbf{G}$, and write $\mathbf{M}_{a}$ for the maximal $\mathbb{Q}$-anisotropic subgroup in $\mathbf{M}_{0}$ so that $\mathbf{M}_{0}=\mathbf{T}\cdot\mathbf{M}_{a}$.
Now we choose a maximal $\mathbb{Q}$-torus $\mathbf{S}$ in $\mathbf{P}_{0}$ containing $\mathbf{T}$. The Lie algebra $\mathfrak{g}$ of $\mathbf{G}$ can be written as a direct sum
$\mathfrak{g}=\mathfrak{g}_{0}\oplus\bigoplus_{\alpha\in\Psi}\mathfrak{g}_{\alpha}$
of root spaces relative to $\mathbf{S}$ via the adjoint representation of $\mathbf{G}$, where $\Psi$ denotes the set of roots in $\mathfrak{g}$ relative to $\mathbf{S}$. We may determine a set of positive roots in $\Psi$, which we denote by $\Psi^{+}$, such that
$\operatorname{Lie}(R_{u}(\mathbf{P}_{0}))\subset\sum_{\alpha\in\Psi^{+}}\mathfrak{g}_{\alpha}\textup{ and }\operatorname{Lie}(R_{u}(\overline{\mathbf{P}}_{0}))\subset\sum_{\alpha\in\Psi\setminus\Psi^{+}}\mathfrak{g}_{\alpha}.$
The set of simple roots in $\Psi^{+}$ is denoted by $\Pi$.
It is known that there are complete classifications of irreducible representations of complex semisimple groups and Lie algebras [21, 27]. Let $V=\bigoplus_{\lambda}V_{\lambda}$ be the decomposition of the vector space $V$ in $\rho$ into the direct sum of weight spaces $V_{\lambda}$ relative to $\mathbf{S}$. According to the theorem of highest weight, there is a unique highest weight $\lambda_{0}$ among the weights $\lambda$’s (the order is determined by $(\Psi,\Psi^{+},\Pi)$) such that
  1. (1)
    $\dim_{\mathbb{C}}V_{\lambda_{0}}=1$.
  2. (2)
    for any $\alpha\in\Psi^{+}$, any $E_{\alpha}\in\mathfrak{g}_{\alpha}$ annihilates $V_{\lambda_{0}}$ via the differential $d\rho$ of $\rho$, any $n_{\alpha}\in\exp(\mathfrak{g}_{\alpha})$ fixes elements in $V_{\lambda_{0}}$ via $\rho$ (where $\exp$ is the exponential map), and elements of $V_{\lambda_{0}}$ are the only vectors with this property.
  3. (3)
    every weight $\lambda$ in $\rho$ is of the form $\lambda_{0}-\sum_{i=1}^{l}n_{i}\alpha_{i}$ where $n_{i}\in\mathbb{Z}_{\geq 0}$ and $\alpha_{i}\in\Pi$.
Note that by our choices of $\Phi$ and $\Psi$, we have
$\{\alpha|_{\mathbf{T}}:\alpha\in\Psi,\alpha|_{\mathbf{T}}\neq 0\}=\{\alpha:\alpha\in\Phi\}.$
On the other hand, one can also write $V$ as a direct sum of weight spaces relative to $\mathbf{T}$ as follows
$V=\bigoplus_{\beta}V_{\beta}.$
Let $\beta_{0}$ be the weight in the decomposition above such that its weight space $V_{\beta_{0}}$ contains $V_{\lambda_{0}}$. Note that $\beta_{0}$ is defined over $\mathbb{Q}$ and $\beta_{0}=\lambda_{0}|_{\mathbf{T}}$. In particular, $\lambda_{0}(\mathbf{T}(\mathbb{R}))\subset\mathbb{R}$ and $\beta_{0}$ is the highest weight among $\beta$’s relative to $\mathbf{T}$ (the order is determined by the root system $(\Phi,\Phi^{+},\Delta)$) (as discussed in §1).
Since the stable subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in $R_{u}(\mathbf{P}_{0})$, we have $\alpha(a_{t})\leq 0$ $(t\geq 0)$ for any $\alpha\in\Phi^{+}$. From property (3) of the representation $\rho$ above and the fact that $\textup{Im}\rho\subset\operatorname{\mathbf{SL}}(V)$, one can deduce that $\beta_{0}(a_{t})=\lambda_{0}(a_{t})\leq 0$. If $\beta_{0}(a_{t})=\lambda_{0}(a_{t})=0$, then by the fact that $\textup{Im}\rho\subset\operatorname{\mathbf{SL}}(V)$, for any weight $\lambda$ in $\rho$ relative to $\mathbf{S}$, we have $\lambda(a_{t})=0$ and $\{a_{t}\}_{t\in\mathbb{R}}\subset\ker\rho$, which contradicts the assumption that $\dim\ker\rho=0$. Therefore we have $\beta_{0}(a_{t})=\lambda_{0}(a_{t})<0$ and
$\rho(a_{t})\cdot v=e^{\beta_{0}(a_{t})}\cdot v\to 0\textup{ as }t\to\infty$
for any $v\in V_{\beta_{0}}$ (here $\beta_{0}(a_{-1})>0$ is the fastest contracting rate under the action of $\{a_{t}\}_{t\in\mathbb{R}}$).
Let $\mathbf{P}_{\beta_{0}}$ be the stabilizer of the weight space $V_{\beta_{0}}$ in $\mathbf{G}$. Then $\mathbf{P}_{\beta_{0}}$ contains the minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$ since $\beta_{0}$ is the highest weight among the weights $\beta$’s relative to $\mathbf{T}$ in $\rho$. Therefore, $\mathbf{P}_{\beta_{0}}$ is a parabolic $\mathbb{Q}$-subgroup in $\mathbf{G}$. Let $\overline{\mathbf{P}}_{\beta_{0}}$ be the opposite parabolic $\mathbb{Q}$-subgroup of $\mathbf{P}_{\beta_{0}}$ containing $\overline{\mathbf{P}}_{0}$, and denote by $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ the unipotent radical of $\overline{\mathbf{P}}_{\beta_{0}}$.
Let $N(\mathbf{T})$ and $Z(\mathbf{T})$ be the normalizer and centralizer of $\mathbf{T}$ respectively. Then the Weyl group relative to $\mathbb{Q}$ is defined by
${}_{\mathbb{Q}}{W}=N(\mathbf{T})/Z(\mathbf{T}).$
Let $\mathbf{P}_{0}$ be the minimal parabolic $\mathbb{Q}$-subgroup of $\mathbf{G}$ as defined above. Then the Bruhat decomposition of $\mathbf{G}$ is the following [6, §21]
$\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot\mathbf{P}_{0}(\mathbb{Q})$
which implies that
$\quad\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot\overline{\mathbf{P}}_{0}(\mathbb{Q}),\quad\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{Q}).$
If we choose a set $\{w_{i}\}_{i\in I}$ of representatives of ${}_{\mathbb{Q}}{W}$ in $\mathbf{G}(\mathbb{Q})$, then for each $w_{i}$, define
$\mathbf{F}_{w_{i}}=R_{u}(\overline{\mathbf{P}}_{0})\cap w_{i}^{-1}R_{u}(\overline{\mathbf{P}}_{\beta_{0}})w_{i},\quad\mathbf{H}_{w_{i}}=R_{u}(\overline{\mathbf{P}}_{0})\cap w_{i}^{-1}\mathbf{P}_{\beta_{0}}w_{i}$
and we have
$R_{u}(\overline{\mathbf{P}}_{0})=\mathbf{H}_{w_{i}}\cdot\mathbf{F}_{w_{i}}.$
Then one can further write the Bruhat decomposition as
$\mathbf{G}(\mathbb{Q})=$ $\bigcup_{i\in I}\mathbf{P}_{0}(\mathbb{Q})\cdot(w_{i}\mathbf{H}_{w_{i}}w_{i}^{-1})(\mathbb{Q})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{Q}).$
$=$ $\bigcup_{i\in\bar{I}}\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{Q})$
where the index sets $I$ and $\bar{I}$ are defined as in §1. We denote by
$\mathcal{C}_{\bar{I}}:=\bigcup_{i\in\bar{I}}w_{i}\mathbf{F}_{w_{i}}(\mathbb{R}).$
Note that the group $w_{i}\mathbf{H}_{w_{i}}w_{i}^{-1}\subset\mathbf{P}_{\beta_{0}}$ is generated by unipotent subgroups whose Lie algebras are direct sums of $\mathbb{Q}$-root spaces in $\mathbf{G}$ relative to $\mathbf{T}$, and $w_{i}\mathbf{H}_{w_{i}}w_{i}^{-1}$ fixes every element in $V_{\beta_{0}}$ by the structure of the representation $\rho$.
Definition 2.1.
An element $g\in\mathbf{G}(\mathbb{R})$ is called rational if $\rho(g)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}}$ is Zariski dense in $V_{\beta_{0}}$.
Lemma 2.2.
An element $g\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})$ is rational if and only if $g\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q})$.
Proof.
Let $g\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})$ be a rational element in $\mathbf{G}(\mathbb{R})$. Then by definition, there exists a discrete subgroup $\Lambda_{g}\subset\mathbb{Z}^{d}$ such that $\rho(g)\cdot\Lambda_{g}$ is a lattice in $V_{\beta_{0}}(\mathbb{R})$. Choose $\sigma\in\operatorname{Gal}(\mathbb{R}/\mathbb{Q})$. Since $V_{\beta_{0}}$ is defined over $\mathbb{Q}$, we have
$\sigma(\rho(g)\cdot\Lambda_{g})=\rho(\sigma(g))\cdot\Lambda_{g}\subset\sigma(V_{\beta_{0}}(\mathbb{R}))=V_{\beta_{0}}(\mathbb{R})$
and $\rho(\sigma(g))\cdot\Lambda_{g}$ is Zariski dense in $V_{\beta_{0}}$. Then we obtain that
$\rho(\sigma(g)g^{-1})V_{\beta_{0}}=\rho(\sigma(g)g^{-1})\overline{\rho(g)\cdot\Lambda_{g}}=V_{\beta_{0}}$
and $\sigma(g)g^{-1}$ is in the stabilizer $\mathbf{P}_{\beta_{0}}$ of $V_{\beta_{0}}$. Here $\overline{\rho(g)\cdot\Lambda_{g}}$ denotes the Zariski closure of $\rho(g)\cdot\Lambda_{g}$. On the other hand, $\sigma(g)g^{-1}\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ and $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})\cap\mathbf{P}_{\beta_{0}}=\{e\}$. Therefore, $\sigma(g)=g$ for any $\sigma\in\operatorname{Gal}(\mathbb{R}/\mathbb{Q})$, and $g\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q})$. The other direction is clear. ∎
Corollary 2.3.
Let $i\in I$ and $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Then $w_{i}\cdot g$ is rational if and only if $g\in\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{F}_{w_{i}}(\mathbb{Q})$.
Proof.
Let $h_{i}\in\mathbf{H}_{w_{i}}(\mathbb{R})$ and $f_{i}\in\mathbf{F}_{w_{i}}(\mathbb{R})$ such that $g=h_{i}\cdot f_{i}$. Suppose that $w_{i}\cdot g$ is rational. Since $w_{i}\in\mathbf{G}(\mathbb{Q})$, by definition, $w_{i}\cdot h_{i}\cdot f_{i}\cdot w_{i}^{-1}$ is also rational in $\mathbf{G}(\mathbb{R})$. We know that $w_{i}\cdot h_{i}\cdot w_{i}^{-1}$ preserves $V_{\beta_{0}}$, so $w_{i}\cdot f_{i}\cdot w_{i}^{-1}$ is rational. By the fact that $w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}\subset R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ and Lemma 2.2, we conclude that $w_{i}\cdot f_{i}\cdot w_{i}^{-1}\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q})$ and $f_{i}\in\mathbf{F}_{w_{i}}(\mathbb{Q})$.
Conversely, suppose that $g=h_{i}\cdot f_{i}$ where $h_{i}\in\mathbf{H}_{w_{i}}(\mathbb{R})$ and $f_{i}\in\mathbf{F}_{w_{i}}(\mathbb{Q})$. Then
$w_{i}\cdot g=(w_{i}h_{i}w_{i}^{-1})\cdot(w_{i}f_{i}w_{i}^{-1})\cdot w_{i}.$
Note that $w_{i}h_{i}w_{i}^{-1}\in\mathbf{P}_{\beta_{0}}$ and $w_{i}\in\mathbf{G}(\mathbb{Q})$. By definition, we know that $w_{i}\cdot g$ is rational. This completes the proof of the lemma. ∎
By properties of the subset $\mathcal{K}\subset\mathbf{G}(\mathbb{Q})$ in the reduction theory, we can give another characterization of rational elements in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
Lemma 2.4.
Let $i\in I$ and $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Then $w_{i}\cdot g$ is rational if and only if
$g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\cap(\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{R})\cdot\mathcal{K}\cdot\Gamma).$
Proof.
If $w_{i}\cdot g$ is rational, then by Corollary 2.3
$g\in\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{F}_{w_{i}}(\mathbb{Q})\subset\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{G}(\mathbb{Q})=\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{Q})\cdot\mathcal{K}\cdot\Gamma\subset\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{R})\cdot\mathcal{K}\cdot\Gamma.$
Suppose that $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\cap(\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{R})\cdot\mathcal{K}\cdot\Gamma)$. Then there exist $h\in\mathbf{H}_{w_{i}}(\mathbb{R})$ and $f\in\mathbf{F}_{w_{i}}(\mathbb{R})$ such that
$g=h\cdot f,\quad f\in\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{R})\cdot\mathcal{K}\cdot\Gamma.$
We write
$f=h_{i}\cdot p\cdot x\cdot\gamma$
for some $h_{i}\in\mathbf{H}_{w_{i}}(\mathbb{R})$, $p\in\mathbf{P}_{0}(\mathbb{R})$, $x\in\mathcal{K}$ and $\gamma\in\Gamma$. Let $\sigma\in\operatorname{Gal}(\mathbb{R}/\mathbb{Q})$. Then we have
$\sigma(f)=\sigma(h_{i})\cdot\sigma(p)\cdot x\cdot\gamma.$
This implies that $x\cdot\gamma=p^{-1}\cdot h_{i}^{-1}\cdot f=\sigma(p)^{-1}\cdot\sigma(h_{i})^{-1}\cdot\sigma(f)$. Since the product map
$\mathbf{P}_{0}\times\mathbf{H}_{w_{i}}\times\mathbf{F}_{w_{i}}\to\mathbf{G}$
is injective, we have $\sigma(f)=f$ for any $\sigma\in\operatorname{Gal}(\mathbb{R}/\mathbb{Q})$ and hence $f\in\mathbf{F}_{w_{i}}(\mathbb{Q})$. By Corollary 2.3, $w_{i}\cdot g$ is rational. ∎
Definition 2.5.
Let $g$ be a rational element in $\mathbf{G}(\mathbb{R})$. The denominator (or the height) $d(g)$ of $g$ is defined to be the co-volume of the lattice $\rho(g)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}}$ in $V_{\beta_{0}}$.
In the following lemma, we discuss the relation between the shortest vector and the co-volume of a discrete subgroup of the form $\rho(g)\cdot\mathbb{Z}^{n}\cap V_{\beta_{0}}$.
Lemma 2.6.
Let $i\in I$ and $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Suppose that $w_{i}\cdot g$ is rational. Then
$d(w_{i}\cdot g)\sim\delta(\rho(w_{i}\cdot g)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}})^{\dim V_{\beta_{0}}}.$
Here the implicit constant depends only on $\mathbf{G}$ and $\Gamma$.
Proof.
By Corollary 2.3, there exist $h_{i}\in\mathbf{H}_{w_{i}}(\mathbb{R})$ and $f_{i}\in\mathbf{F}_{w_{i}}(\mathbb{Q})$ such that $g=h_{i}\cdot f_{i}$. Note that
$f_{i}\in\mathbf{G}(\mathbb{Q})=w_{i}^{-1}\cdot\mathbf{G}(\mathbb{Q})=w_{i}^{-1}\cdot\mathbf{P}_{0}(\mathbb{Q})\cdot\mathcal{K}\cdot\Gamma$
and there exist $p\in\mathbf{P}_{0}(\mathbb{Q})$, $x\in\mathcal{K}$ and $\gamma\in\Gamma$ such that $f_{i}=w_{i}^{-1}\cdot p\cdot x\cdot\gamma$. We know that $\mathbf{P}_{0}=\mathbf{M}_{0}\cdot R_{u}(\mathbf{P}_{0})$ where $\mathbf{M}_{0}$ is the Levi factor of $\mathbf{P}_{0}$. We write $\mathbf{M}_{a}$ for the maximal $\mathbb{Q}$-anisotropic subgroup in $\mathbf{M}_{0}$. Then $\mathbf{M}_{0}=\mathbf{T}\cdot\mathbf{M}_{a}$, $\mathbf{T}$ commutes with $\mathbf{M}_{a}$ and $\mathbf{M}_{a}(\mathbb{R})/(\mathbf{M}_{a}(\mathbb{Z})\cap\Gamma)$ is compact. So there exist $p_{1}\in\mathbf{T}$, $p_{2}$ in a compact fundamental domain of $\mathbf{M}_{a}(\mathbb{Z})\cap\Gamma$ in $\mathbf{M}_{a}(\mathbb{R})$, $p_{3}\in\mathbf{M}_{a}(\mathbb{Z})\cap\Gamma$ and $u\in R_{u}(\mathbf{P}_{0})$ such that
$p=u\cdot p_{1}\cdot p_{2}\cdot p_{3}.$
Then we have
$w_{i}\cdot g=(w_{i}\cdot h_{i}\cdot w_{i}^{-1})\cdot u\cdot p_{1}\cdot p_{2}\cdot p_{3}\cdot x\cdot\gamma.$
Since $w_{i}\mathbf{H}_{w_{i}}w_{i}^{-1}$ and $R_{u}(\mathbf{P}_{0})$ fix every element in $V_{\beta_{0}}$, and $\mathbf{T}$ and $\mathbf{M}_{a}$ preserve the weight space $V_{\beta_{0}}$, we obtain
$\rho(w_{i}\cdot g)\mathbb{Z}^{d}\cap V_{\beta_{0}}=\rho(p_{1}\cdot p_{2}\cdot p_{3}\cdot x)\mathbb{Z}^{d}\cap V_{\beta_{0}}=\rho(p_{1}\cdot p_{2})(\rho(p_{3}\cdot x)\mathbb{Z}^{d}\cap V_{\beta_{0}}).$
The lemma then follows from the facts that $p_{1}\in\mathbf{T}$ acts as scalars in $V_{\beta_{0}}$, $p_{2}$ is in a fixed compact subset in $\mathbf{M}_{a}(\mathbb{R})$ and $\rho(p_{3}\cdot x)\mathbb{Z}^{d}$ is commensurable with $\mathbb{Z}^{d}$ by $\Gamma$ and $\mathcal{K}$. ∎
Now we consider the case that $w_{i}=e$ is the identity element in ${}_{\mathbb{Q}}W$. Let $g$ be a rational element in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Then $\mathbf{F}_{w_{i}}=\mathbf{F}_{e}$ and $\mathbf{H}_{w_{i}}=\mathbf{H}_{e}$. By Lemma 2.4, there exist $h\in\mathbf{H}_{e}(\mathbb{R})$, $p\in\mathbf{P}_{0}(\mathbb{R})$, $x\in\mathcal{K}$ and $\gamma\in\Gamma$ such that
$g=h\cdot p\cdot x\cdot\gamma.$
Furthermore, we can write
$p=a\cdot m\cdot u$
where $a\in\mathbf{T}(\mathbb{R})$, $m\in\mathbf{M}_{a}(\mathbb{R})$ and $u\in R_{u}(\mathbf{P}_{0})(\mathbb{R})$. Then we can compute the denominator $d(g)$ of $g$ as follows:
$(\rho(g)\cdot\mathbb{Z}^{d})\cap V_{\beta_{0}}=\rho(a\cdot m\cdot u\cdot x)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}}=\rho(a\cdot m\cdot x)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}}=\rho(a\cdot m)(\rho(x)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}})$
and
$d(g)=c_{x}\cdot e^{\beta_{0}(a)\cdot\dim V_{\beta_{0}}}$
for some constant $c_{x}>0$ depending only on $x$. Here we use the facts that $\mathbf{M}_{a}$ and $R_{u}(\mathbf{P}_{0})$ stabilize $V_{\beta_{0}}$ and preserve the volumes of sets in $V_{\beta_{0}}$, $\mathbf{T}$ acts as scalars in $V_{\beta_{0}}$ and $x\in\mathbf{G}(\mathbb{Q})$.
Definition 2.7.
Let $\mathcal{K}=\{x_{j}\}_{j\in J}$. A rational element $g$ in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ is called $j$-rational for some $j\in J$ if it can be written as
$g=h\cdot p\cdot x_{j}\cdot\gamma$
for some $h\in\mathbf{H}_{e}(\mathbb{R})$, $p\in\mathbf{P}_{0}(\mathbb{R})$, $x_{j}\in\mathcal{K}$ and $\gamma\in\Gamma$.
3. Counting rational points
In this section, we consider the problem of counting rational points in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. To estimate the Hausdorff dimension, one usually constructs boxes centered at certain rational points and obtains a tree-like subset. At each level of the tree-like subset, the boxes are disjoint. In our case, it may happen that the boxes at each level are not disjoint if we proceed in the usual way. It implies that the boxes constructed are too many, and we have to sieve out some proportion of the rational points so that the boxes centered at the remaining rational points are disjoint. It is also required that the number of the remaining rational points is not small compared to the total number of rational points in order to avoid any loss of Hausdorff dimension. For our purpose, in the following, we will first sieve some proportion of rational points according to several algebraic conditions, and then count the remaining rational points by the mixing property of the flow $\{a_{t}\}_{t\in\mathbb{R}}$. The disjointness property of the boxes centered at the remaining rational points will follow from the transversal structure of some submanifolds in the homogeneous space $\mathbf{G}(\mathbb{R})^{0}/\Gamma$.
For any $\mathbb{Q}$-algebraic group $\mathbf{L}$ in $\mathbf{G}$ and $\delta>0$, we write $B_{\mathbf{L}}(\delta)$ for the open ball of radius $\delta>0$ around the identity in $\mathbf{L}(\mathbb{R})$. For any $t\in\mathbb{R}$, we write
$B_{\mathbf{L}}(\delta,t):=a_{-t}\cdot B_{\mathbf{L}}(\delta)\cdot a_{t}.$
The connected component of identity of $\mathbf{L}$ in the Zariski topology is denoted by $\mathbf{L}^{0}$, and the connected component of identity in the Lie group $\mathbf{L}(\mathbb{R})$ is denoted by $\mathbf{L}(\mathbb{R})^{0}$. We write $A\lesssim B$ $(A\gtrsim B)$ if there exists a constant $c>0$ such that
$A\leq c\cdot B\quad(A\geq c\cdot B).$
If $A\lesssim B$ and $A\gtrsim B$, then we write $A\sim B$. We will specify the implicit constants in the contexts if necessary.
In this section and next section, we assume that the action of the one-parameter subgroup $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$ is mixing. This holds true when the one-parameter subgroup $\{a_{t}\}_{t\in\mathbb{R}}$ projects nontrivially into any $\mathbb{Q}$-simple factor of $\mathbf{G}$. Later in §5, we will explain how to establish the results of counting rational points and Hausdorff dimension estimates in the case that $\{a_{t}\}_{t\in\mathbb{R}}$ appears only in some $\mathbb{Q}$-simple factors of $\mathbf{G}$.
Let $U$ be an open bounded subset in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. For any $0<A<B$, define
$S(U,A,B)=\{q\in U:q\textup{ rational and }A\leq d(q)\leq B\}.$
We know that the Lie algebra $\mathfrak{a}$ of $\mathbf{T}(\mathbb{R})$ has the following decomposition
$\mathfrak{a}=\operatorname{Lie}(a_{t})\oplus\ker(\beta_{0})$
where $\operatorname{Lie}(a_{t})$ is the Lie algebra of the one parameter subgroup $\{a_{t}\}_{t\in\mathbb{R}}$ and we have
$\beta_{0}(a_{t})=\lambda_{0}(a_{t})<0.$
Denote by $\pi_{\ker(\beta_{0})}$ the projection from $\mathfrak{a}$ to $\ker(\beta_{0})$ according to this decomposition. We also write $\pi_{\ker(\beta_{0})}(a)$ for $\pi_{\ker(\beta_{0})}(\log(a))$ whenever $a\in\exp(\mathfrak{a})$. Here $\log$ is the inverse of the exponential map $\exp$. For any compact subset $L$ in $\operatorname{Lie}(a_{t})$ and any compact subset $K$ in $\ker(\beta_{0})$, we denote by
$\mathfrak{a}_{L,K}=\{x\in\mathfrak{a}:x=y_{1}+y_{2},y_{1}\in L,y_{2}\in K\}.$
For any $x_{j}\in\mathcal{K}$, any compact subset $K_{1}$ in $\ker(\beta_{0})$, any compact subset $K_{2}$ in $\mathbf{H}_{e}(\mathbb{R})$, any compact subset $K_{3}\subset\mathbf{M}_{a}(\mathbb{R})\cap\mathbf{G}(\mathbb{R})^{0}$ and any compact subset $K_{4}\subset R_{u}(\mathbf{P}_{0})(\mathbb{R})$, we define
$S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)$
to be the set of all rational points $q$ in $U$ such that
  1. (1)
    $A\leq d(q)\leq B$ and $q$ is $j$-rational for $x_{j}\in\mathcal{K}$;
  2. (2)
    $q=a\cdot h\cdot m\cdot u\cdot x_{j}\cdot\gamma$ for some $a\in\exp(\mathfrak{a})$, $\pi_{\ker(\beta_{0})}(a)\in K_{1}$ and $h\in K_{2}$, $m\in K_{3}$, $u\in K_{4}$ and $\gamma\in\Gamma$.
Note that $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)\subset S(U,A,B)$, and by definition one can check that if
$S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)\neq\emptyset,$
then $x_{j}\in\mathbf{G}(\mathbb{R})^{0}$ (and such elements exist as $e\in\mathcal{K}$). The elements in $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)$ are the rational points we are interested in when we construct a tree-like subset in §4. We denote by
$S_{K_{1},K_{2},K_{3},K_{4}}(U,A,B):=\bigcup_{j\in J}S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B).$
To count the rational points in $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)$, we need results about limiting distributions of translates of unipotent orbits pushed by the one-parameter semisimple flow $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$. The following is a direct consequence of the mixing property of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$.
Proposition 3.1.
Let $x\in\mathbf{G}(\mathbb{R})^{0}/\Gamma$ and $W\subset\mathbf{G}(\mathbb{R})^{0}/\Gamma$ an open bounded subset whose boundary has measure zero with respect to the invariant probability measure $\mu_{\mathbf{G}^{0}(\mathbb{R})/\Gamma}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$. Let $\chi_{W}$ denote the characteristic function of $W$. Then for any bounded open subset $U$ in $R_{u}(\overline{\mathbf{P}}_{0})$ we have
[…]
where $\mu_{R_{u}(\overline{\mathbf{P}}_{0})}$ is the Haar measure on $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
Remark 3.2.
Note that $R_{u}(\overline{\mathbf{P}}_{0})$ is not necessarily the unstable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$. Here we can still apply the mixing property to obtain Proposition 3.1 as long as $R_{u}(\overline{\mathbf{P}}_{0})$ is contained in the group generated by the unstable subgroup and the central subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$. (See [25, §2].)
Now we can follow the same arguments as in [17, 48] and measure the size of the subset $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)$ for a fixed index $j\in J$ with $x_{j}\in\mathbf{G}(\mathbb{R})^{0}$, and for some sufficiently large numbers $B>A>0$. Recall that
$R_{u}(\overline{\mathbf{P}}_{0})=\mathbf{H}_{e}\cdot\mathbf{F}_{e}$
where $\mathbf{F}_{e}=R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ and $\mathbf{H}_{e}=R_{u}(\overline{\mathbf{P}}_{0})\cap\mathbf{P}_{\beta_{0}}$. We fix a Haar measure $\mu_{\mathbf{H}_{e}}$ on $\mathbf{H}_{e}(\mathbb{R})$ and a Haar measure $\mu_{\mathbf{F}_{e}}$ on $\mathbf{F}_{e}(\mathbb{R})$ such that the product maps
$\mathbf{F}_{e}\times\mathbf{H}_{e}\to\mathbf{F}_{e}\cdot\mathbf{H}_{e}=R_{u}(\overline{\mathbf{P}}_{0})\textup{ and }\mathbf{H}_{e}\times\mathbf{F}_{e}\to\mathbf{H}_{e}\cdot\mathbf{F}_{e}=R_{u}(\overline{\mathbf{P}}_{0})$
induce Haar measures on $R_{u}(\overline{\mathbf{P}}_{0}).$ Then for any $q\in\mathbf{F}_{e}(\mathbb{Q})$, we define $m_{\mathbf{H}_{e}q}$ to be the locally finite measure defined on $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ which is supported on $\mathbf{H}_{e}(\mathbb{R})\cdot q$ and induced by $\mu_{\mathbf{H}_{e}}$ via the product map
$\mathbf{H}_{e}\times\{q\}\to\mathbf{H}_{e}\cdot q\subset R_{u}(\overline{\mathbf{P}}_{0}).$
We define
$m_{\mathbf{H}_{e}}:=\sum_{q\in\mathbf{F}_{e}(\mathbb{Q})}m_{\mathbf{H}_{e}q}.$
Note that $m_{\mathbf{H}_{e}}$ is not a locally finite measure on $R_{u}(\overline{\mathbf{P}}_{0})$, and it is defined by the leaf-wise measures on the countable leaves through the rational elements in $\mathbf{F}_{e}(\mathbb{Q})$ in the foliation $\mathcal{F}_{\mathbf{H}_{e}}$ induced by the group action of $\mathbf{H}_{e}$ on $R_{u}(\overline{\mathbf{P}}_{0})$. Note also that by Corollary 2.3, we have
$S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)\subset\bigcup_{q\in\mathbf{F}_{e}(\mathbb{Q})}\mathbf{H}_{e}(\mathbb{R})\cdot q=\mathbf{H}_{e}(\mathbb{R})\cdot\mathbf{F}_{e}(\mathbb{Q}).$
In the following, we estimate the size of the subset
$S_{K_{1},K_{2},K_{3},K_{4},j}(U,(l/2)^{\dim V_{\beta_{0}}},l^{\dim V_{\beta_{0}}})$
with respect to the measure $m_{\mathbf{H}_{e}}$ for sufficiently large $l>0$. For convenience, we write
$A_{l}=(l/2)^{\dim V_{\beta_{0}}}\textup{ and }B_{l}=l^{\dim V_{\beta_{0}}}\quad(\forall l>0).$
For any $l>1$, let $T=T(l)>0$ such that
$\beta_{0}(a_{T})=-\ln l.$
Let $q$ be a rational element in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. By Lemma 2.4, we may write
$q=a\cdot h\cdot m\cdot u\cdot x_{k}\cdot\gamma\in\mathbf{H}_{e}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{R})\cdot x_{k}\cdot\Gamma$
for some $a\in\mathbf{T}(\mathbb{R})$, $h\in\mathbf{H}_{e}(\mathbb{R})$, $m\in\mathbf{M}_{a}(\mathbb{R})$, $u\in R_{u}(\mathbf{P}_{0})(\mathbb{R})$, $x_{k}\in C$ and $\gamma\in\Gamma$. Then
$q\in S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$
$\iff$ $q\in U,\;A_{l}\leq d(q)\leq B_{l},\textrm{ and }a\in\exp(\mathfrak{a}),\pi_{\ker(\beta_{0})}(a)\in K_{1},h\in K_{2},m\in K_{3},u\in K_{4},k=j.$
$\iff$ $a_{T}\cdot q\Gamma\in a_{T}\cdot U\Gamma/\Gamma\textrm{ and }a_{T}\cdot q\Gamma\in\exp(\mathfrak{a}_{I_{0},K_{1}})\cdot K_{2}\cdot K_{3}\cdot K_{4}\cdot x_{j}\cdot\Gamma/\Gamma$
where $I_{0}$ is the following compact interval in the Lie algebra of $\left\{a_{t}\right\}_{t\in\mathbb{R}}$
$I_{0}:=\left\{x\in\mathbb{\operatorname{Lie}}(a_{t}):-\ln(2/{c_{x_{j}}^{\frac{1}{\dim V_{\beta_{0}}}}})\leq\beta_{0}(x)\leq-\ln(1/{c_{x_{j}}^{\frac{1}{\dim V_{\beta_{0}}}}})\right\}.$
Here we use the formula
$d(q)=c_{x_{j}}e^{\beta_{0}(a)\cdot\dim V_{\beta_{0}}}.$
This implies that
$a_{T}\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma=a_{T}\cdot U\Gamma/\Gamma\cap\exp(\mathfrak{a}_{I_{0},K_{1}})K_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma.$
Since $\exp(\mathfrak{a}_{I_{0},K_{1}})K_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma$ is a compact subset in $\mathbf{G}(\mathbb{R})^{0}/\Gamma$, there exist $\delta_{0}>0$ and a small neighborhood of identity $B_{\mathbf{F}_{e}}(\delta_{0})$ in $\mathbf{F}_{e}(\mathbb{R})$ such that
$B_{\mathbf{F}_{e}}(\delta_{0})\times\exp(\mathfrak{a}_{I_{0},K_{1}})K_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma\to B_{\mathbf{F}_{e}}(\delta_{0})\exp(\mathfrak{a}_{I_{0},K_{1}})K_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma$
is a homeomorphism. We conclude that the following product map
$B_{\mathbf{F}_{e}}(\delta_{0})\times a_{T}\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma\to B_{\mathbf{F}_{e}}(\delta_{0})\cdot a_{T}\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma$
gives a transversal structure near the set $a_{T}\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma$ in $\mathbf{G}(\mathbb{R})^{0}/\Gamma$. Consequently, for any $p,q\in S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$, the subsets $B_{\mathbf{F}_{e}}(\delta_{0},T)\cdot p\Gamma$ and $B_{\mathbf{F}_{e}}(\delta_{0},T)\cdot q\Gamma$ are disjoint, where
$B_{\mathbf{F}_{e}}(\delta_{0},T):=a_{-T}\cdot B_{\mathbf{F}_{e}}(\delta_{0})\cdot a_{T}.$
Now we estimate the size of $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$ with respect to the measure $m_{\mathbf{H}_{e}}$. First we prove an upper bound for $m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))$. Fix a sufficiently small number $0<\epsilon<\delta_{0}$ such that
$\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U)\leq\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(B_{\mathbf{F}_{e}}(\epsilon)\cdot U)\leq 2\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U).$
Then for sufficiently large $l>0$, we have
$B_{\mathbf{F}_{e}}(\epsilon,T)\subset B_{\mathbf{F}_{e}}(\epsilon)\textup{ and}B_{\mathbf{F}_{e}}(\epsilon,T)\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma\subset B_{\mathbf{F}_{e}}(\epsilon)\cdot U\Gamma/\Gamma.$
Since $B_{\mathbf{F}_{e}}(\delta_{0},T)$ and $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$ form a transversal structure near $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$, by Proposition 3.1, we deduce that
$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))\mu_{\mathbf{F}_{e}}(B_{\mathbf{F}_{e}}(\epsilon,T))$
[…]
$\sim$ $\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(B_{\mathbf{F}_{e}}(\epsilon)\cdot U)\cdot\mu_{\mathbf{G}(\mathbb{R})^{0}/\Gamma}(B_{\mathbf{F}_{e}}(\epsilon)\exp(\mathfrak{a}_{K_{1},I_{0}})K_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma)$
$\sim$ $\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U)\cdot\mu_{\mathbf{F}_{e}}(B_{\mathbf{F}_{e}}(\epsilon))$
as $T\to\infty$. Here the implicit constant in the last equation depends only on the parameter $\delta_{0}>0$ (as $0<\epsilon<\delta_{0}$), and hence depends only on the compact subsets $K_{i}$’s. Note that
$\mu_{\mathbf{F}_{e}}(B_{\mathbf{F}_{e}}(\epsilon,T))=$ $\mu_{\mathbf{F}_{e}}(B_{\mathbf{F}_{e}}(\epsilon))\cdot e^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{T})}$
$=$ $\mu_{\mathbf{F}_{e}}(B_{\mathbf{F}_{e}}(\epsilon))\cdot l^{\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{T})/\beta_{0}(a_{T})}$
$=$ $\mu_{\mathbf{F}_{e}}(B_{\mathbf{F}_{e}}(\epsilon))\cdot l^{\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}.$
So for sufficiently large $l>0$ we have
$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))\lesssim l^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U)$
where the implicit constant depends only on $K_{1}$, $K_{2}$, $K_{3}$, $K_{4}$, $I_{0}$, $\mathbf{G}$ and $\Gamma$. Similarly, using the arguments in [17, §4], we can prove a lower bound for $m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))$ and for any sufficiently large $l>0$ we have
$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))\gtrsim l^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U).$
To sum up, we have the following
Proposition 3.3.
Let $U$ be a small open bounded subset in $R_{u}(\overline{\mathbf{P}}_{0})$. Then for any $j\in J$ with $x_{j}\in\mathcal{K}\cap\mathbf{G}(\mathbb{R})^{0}$ and any sufficiently large $l>0$, we have
$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))\sim l^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U)$
and
$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4}}(U,A_{l},B_{l}))\sim l^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U).$
Here the implicit constants depend only on the compact subsets $K_{i}$’s, $\mathbf{G}$ and $\Gamma$.
Remark 3.4.
For our purpose, in the following, we fix some compact subsets $K_{1}\subset\ker\beta_{0}$, $K_{2}\subset\mathbf{H}_{e}(\mathbb{R})$, $K_{3}\subset\mathbf{M}_{a}(\mathbb{R})\cap\mathbf{G}(\mathbb{R})^{0}$ and $K_{4}\subset R_{u}(\mathbf{P}_{0})(\mathbb{R})$ discussed above.
4. A lower bound for the Hausdorff dimension of $S_{\gamma}^{c}$
In this section, we prove Theorem 1.3 under the assumption that the action of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$ is mixing.
Lemma 4.1.
Fix $j\in J$ with $x_{j}\in\mathbf{G}(\mathbb{R})^{0}$. Let $U$ be an open bounded subset in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ and
$\beta_{0}(a_{T})=-\ln l$
for some $T>0$ and $l>1$. Let $\mathcal{F}_{q}=\mathbf{H}_{e}(\mathbb{R})\cdot q$ be the leaf through $q\in\mathbf{F}_{e}(\mathbb{Q})$ such that
$\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\neq\emptyset.$
Then there exist $\theta_{1}>0$ and $\theta_{2}>0$ such that for any $p\in\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$
$B_{\mathbf{H}_{e}}(\theta_{1},T)\cdot p\cap U\subset\mathcal{F}_{q}\cap S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$
where $\tilde{K}_{2}=B_{\mathbf{H}_{e}}(\theta_{2})\cdot K_{2}$ and $B_{\mathbf{H}_{e}}(\theta_{1},T)=a_{-T}\cdot B_{\mathbf{H}_{e}}(\theta_{1})\cdot a_{T}$. Here the constants $\theta_{1}$ and $\theta_{2}$ depend only on $K_{i}$’s, $\mathbf{G}$ and $\Gamma$.
Proof.
By the discussion in §3, we know that
$a_{T}\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma=a_{T}\cdot U\Gamma/\Gamma\cap\exp(\mathfrak{a}_{K_{1},I_{0}})K_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma.$
Choose $\theta_{1},\theta_{2}>0$ sufficiently small so that for any $a\in\exp(\mathfrak{a}_{I_{0},K_{1}})$ we have
$a^{-1}B_{\mathbf{H}_{e}}(\theta_{1})a\subset B_{\mathbf{H}_{e}}(\theta_{2}).$
Now for any $p\in\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$, there exist $a\in\exp(\mathfrak{a}_{I_{0},K_{1}})$, $h\in K_{2}$, $m\in K_{3}$, $u\in K_{4}$ and $\gamma\in\Gamma$ such that
$a_{T}\cdot p=a\cdot h\cdot m\cdot u\cdot x_{j}\cdot\gamma.$
Then
$a_{T}\cdot B_{\mathbf{H}_{e}}(\theta_{1},T)\cdot p=B_{\mathbf{H}_{e}}(\theta_{1})\cdot a\cdot h\cdot m\cdot u\cdot\gamma\subset\exp(\mathfrak{a}_{I_{0},K_{1}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma$
where $\tilde{K}_{2}=B_{\mathbf{H}_{e}}(\theta_{2})\cdot K_{2}$. By definition, we have
$B_{\mathbf{H}_{e}}(\theta_{1},T)\cdot p\cap U\subset S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(U,A_{l},B_{l}).$
This completes the proof of the lemma. ∎
Lemma 4.2.
We have
$\max\{\alpha(a_{1}):\alpha\in\Phi(\mathbf{F}_{e})\}>0.$
Proof.
Suppose on the contrary that for all $\alpha\in\Phi(\mathbf{F}_{e})$, $\alpha(a_{1})=0$. Let $\Omega_{\mathbf{F}_{e}}$ and $\Omega_{\mathbf{H}_{e}}$ be small open neighborhoods of identity in $\mathbf{F}_{e}(\mathbb{R})$ and $\mathbf{H}_{e}(\mathbb{R})$ respectively, and
$\Omega_{R_{u}(\overline{\mathbf{P}}_{0})}:=\Omega_{\mathbf{F}_{e}}\cdot\Omega_{\mathbf{H}_{e}}.$
Note that $\mathbf{H}_{e}(\mathbb{R})$ stabilizes every element in $V_{\beta_{0}}$. Now for any $p=f\cdot h\in\Omega_{R_{u}(\overline{\mathbf{P}}_{0})}$ with $f\in\Omega_{\mathbf{F}_{e}}$ and $h\in\Omega_{\mathbf{H}_{e}}$, we have
$a_{t}\cdot p=(a_{t}fa_{-t})(a_{t}ha_{-t})a_{t}=f\cdot(a_{t}ha_{-t})\cdot a_{t}.$
The element $a_{t}$ is rational and
$\delta(\rho(a_{t})\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}})=e^{\beta_{0}(a_{t})}.$
This implies that
$\delta(\rho(a_{t}\cdot p)\cdot\mathbb{Z}^{d})\leq\kappa\cdot e^{-\beta_{0}(a_{-1})t}$
for any $t>0$ where the constant $\kappa$ depends only on $\Omega_{\mathbf{F}_{e}}$. Let $\Omega_{\mathbf{P}_{0}}$ be a small neighborhood of identity in $\mathbf{P}_{0}$. Then one can deduce from the inequality above that for any point $p\Gamma\in\Omega_{\mathbf{P}_{0}}\cdot\Omega_{R_{u}(\overline{\mathbf{P}}_{0})}\Gamma$, the orbit $a_{t}\cdot p\Gamma$ diverges in $\mathbf{G}(\mathbb{R})^{0}/\Gamma$. This contradicts the mixing property of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$. ∎
Let $X$ be a Riemannian manifold, $m$ a volume form on $X$ and $E$ a compact subset of $X$. Denote by $\operatorname{diam}(S)$ the diameter of a set $S\subset X$. A collection $\mathcal{A}$ of compact subsets of $E$ is said to be tree-like if $\mathcal{A}$ is the union of finite sub-collections $\mathcal{A}_{k}$ such that
  1. (1)
    $\mathcal{A}_{0}=\left\{E\right\}$;
  2. (2)
    For any $k$ and $S_{1},S_{2}\in\mathcal{A}_{j}$, either $S_{1}=S_{2}$ or $S_{1}\cap S_{2}=\emptyset$;
  3. (3)
    For any $k$ and $S_{1}\in\mathcal{A}_{k+1}$, there exists $S_{2}\in\mathcal{A}_{k}$ such that $S_{1}\subset S_{2}$;
  4. (4)
    $d_{k}(\mathcal{A}):=\sup_{S\in\mathcal{A}_{k}}\operatorname{diam}(S)\to 0$ as $k\to\infty$.
We write $\mathbf{A}_{k}:=\bigcup_{A\in\mathcal{A}_{k}}A$ and $\mathbf{A}_{\infty}:=\bigcap_{k\in\mathbb{N}}\mathbf{A}_{k}$. We also define
$\Delta_{k}(\mathcal{A}):=\inf_{S\in\mathcal{A}_{k}}\frac{m(\mathbf{A}_{k+1}\cap S)}{m(S)}.$
Theorem 4.3 ([25, 31, 44]).
Let $(X,m)$ be a Riemannian manifold, where $m$ is the volume form on $X$. Then for any tree-like collection $\mathcal{A}$ of subsets of $E$
$\dim_{H}(\mathbf{A}_{\infty})\geq\dim_{H}X-\limsup_{k\to\infty}\frac{\sum_{i=0}^{k}\log(\Delta_{i}(\mathcal{A}))}{\log(d_{k+1}(\mathcal{A}))}.$
Proof of Theorem 1.3.
We fix the compact subsets $K_{1}\subset\ker(\beta_{0}),K_{2}\subset\mathbf{H}_{e}(\mathbb{R}),K_{3}\subset\mathbf{M}_{a}(\mathbb{R})\cap\mathbf{G}(\mathbb{R})^{0},K_{4}\subset R_{u}(\mathbf{P}_{0})$ as in Remark 3.4, and also fix $j\in J$ with $x_{j}\in\mathcal{K}\cap\mathbf{G}(\mathbb{R})^{0}$. Let $\epsilon>0$ be a sufficiently small number. Let $\nu_{0}$ be the $\mathbb{Q}$-root in $R_{u}(\overline{\mathbf{P}}_{0})$ such that
$\nu_{0}(a_{1})=\max\{\alpha(a_{1}):\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{0}))\}.$
Then by Lemma 4.2, $\nu_{0}(a_{1})>0$.
We start with a small open bounded box $U_{0}$ in $R_{u}(\overline{\mathbf{P}}_{0})$.
For $k=0$, we set $\mathcal{A}_{0}=\left\{U_{0}\right\}$.
For $k=1$, we choose sufficiently large numbers $l_{1}>0$ and $T_{1}>0$ such that
$\beta_{0}(a_{T_{1}})=-\ln l_{1}.$
By Proposition 3.3, we know that
$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}}))\sim l_{1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U_{0})$
and
$m_{\mathbf{H}_{e}}(S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}}))\sim l_{1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U_{0})$
where $\tilde{K}_{2}=B_{\mathbf{H}_{e}}(\theta_{2})\cdot K_{2}$ for some $\theta_{2}>0$ as defined in Lemma 4.1. Now for any leaf $\mathcal{F}_{q}=\mathbf{H}_{e}(\mathbb{R})\cdot q$ ($q\in\mathbf{F}_{e}(\mathbb{Q})$) in the foliation $\mathcal{F}_{\mathbf{H}_{e}}$ such that
$\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})\neq\emptyset,$
we devide the region $\mathcal{F}_{q}\cap U_{0}$ into small cubes of side length
$\theta_{1}\cdot\exp(-\nu_{0}(a_{T_{1}}))/10$
where $\theta_{1}>0$ is the constant defined in Lemma 4.1. Then we collect those cubes $R$ which intersect $S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})$, and denote the corresponding collection by
$\mathcal{G}_{1,q}=\{R:R\cap S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})\neq\emptyset\}.$
Note that $\theta_{1}\cdot\exp(-\nu_{0}(a_{T_{1}}))/10$ is smaller than the minimum side length of the rectangle $B_{\mathbf{H}_{e}}(\theta_{1},T_{1})$, so by Lemma 4.1, we know that each cube in $\mathcal{G}_{1,q}$ is contained in
$S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})$
where $\tilde{K}_{2}=B_{\mathbf{H}_{e}}(\theta_{2})\cdot K_{2}$. Let
$\mathcal{H}_{1}=\bigcup_{\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})\neq\emptyset}\bigcup_{R\in\mathcal{G}_{1,q}}R.$
Then we have
$l_{1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U_{0})$
$\sim m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}}))$
$\leq m_{\mathbf{H}_{e}}(\bigcup_{\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})\neq\emptyset}\bigcup_{R\in\mathcal{G}_{1,q}}R)$
$\leq m_{\mathbf{H}_{e}}(S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}}))$
$\sim l_{1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U_{0})$
and
$m_{\mathbf{H}_{e}}(\mathcal{H}_{1})\sim l_{1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U_{0}).$
Note that each cube $R$ in $\mathcal{H}_{1}$ is contained in $S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})$. Let
$t_{1}=\frac{\ln l_{1}}{\beta_{0}(a_{-1})-(\tau+\epsilon)}.$
By the computations in §3, we can choose a sufficiently small number $\tilde{\delta}_{0}>0$ such that
$B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\times\exp(\mathfrak{a}_{K_{1},I_{0}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma\to B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\exp(\mathfrak{a}_{K_{1},I_{0}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma$
is a homeomorphism, and the subsets in the collection
$\mathcal{F}_{1}(U_{0}):=\left\{\left(a_{-t_{1}}\cdot B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\cdot a_{t_{1}}\right)\cdot q:q\in\mathcal{H}_{1}\right\}$
are disjoint by this homeomorphism as
$a_{-t_{1}}\cdot B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\cdot a_{t_{1}}\subset B_{\mathbf{F}_{e}}(\tilde{\delta}_{0},T_{1}).$
We write
$\Phi(\mathbf{F}_{e})=\Phi^{0}(\mathbf{F}_{e})\cup\Phi^{1}(\mathbf{F}_{e})$
where
$\Phi^{0}(\mathbf{F}_{e})=\{\alpha\in\Phi(\mathbf{F}_{e}):\alpha(a_{1})=0\}\textup{ and }\Phi^{1}(\mathbf{F}_{e})=\{\alpha\in\Phi(\mathbf{F}_{e}):\alpha(a_{1})\neq 0\}.$
We denote by
$\mathcal{P}_{1}=\bigcup_{E\in\mathcal{F}_{1}(U_{0})}E.$
Now we can divide the subset $\mathcal{P}_{1}$ into cubes of side length
$\tilde{\delta}_{0}\cdot l_{1}^{\frac{\nu_{0}(a_{-1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}}$
which is smaller than $\theta_{1}\cdot\exp(-\nu_{0}(a_{T_{1}}))/10$ and the minimum side length of the rectangle
$a_{-t_{1}}\cdot B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\cdot a_{t_{1}}$
if $T_{1}>0$ and $l_{1}>0$ are sufficiently large. Note that the set $\Phi^{0}(\mathbf{F}_{e})$ may not be empty and the diameter of the set
$a_{-t_{1}}\cdot B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\cdot a_{t_{1}}$
may be larger than the diameter of $U_{0}$, so some cubes we obtain here from dividing the subset $\mathcal{P}_{1}$ may be outside the set $U_{0}$. For our purpose, we collect only those cubes which are inside the subset $U_{0}$. In this manner, we obtain a family of disjoint cubes constructed from sets in $\mathcal{F}_{1}(U_{0})$ inside $U_{0}$, which we denote by $\mathcal{A}_{1}$.
We remark here that if $\Phi^{0}(\mathbf{F}_{e})=\emptyset$, then all the cubes we obtain from dividing $\mathcal{P}_{1}$ are inside the subset $U_{0}$ if $T_{1}>0$ and $l_{1}>0$ are chosen to be sufficiently large (and also if we shrink or enlarge $U_{0}$ slightly to avoid the complexity caused by the boundary of $U_{0}$). We will see later that in the computations there is little difference between the case $\Phi^{0}(\mathbf{F}_{e})\neq\emptyset$ and the case $\Phi^{0}(\mathbf{F}_{e})=\emptyset$. Indeed, when we apply Theorem 4.3 and calculate $\Delta_{i}(\mathcal{A})$ and $d_{i}(\mathcal{A})$, the difference between these two cases may affect the value of the formula in finite steps, but eventually when we take the limit, this difference will disappear since we keep choosing sufficiently large numbers $T_{1},l_{1}$ and later $T_{2},l_{2},T_{3},l_{3},\dots$ to offset the effects by the difference at early stages.
Similarly, we can construct $\mathcal{A}_{k}$ inductively for any $k\in\mathbb{N}$. For $k>1$ we choose sufficiently large numbers $l_{k}>0$ and $T_{k}>0$ such that $\beta_{0}(a_{\tau_{k}})=-\ln l_{k}$. For each cube $S\in\mathcal{A}_{k-1}$, by the proposition above, we know that
$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}}))\sim l_{k}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(S)$
and
$m_{\mathbf{H}_{e}}(S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}}))\sim l_{k}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(S)$
where $\tilde{K}_{2}=B_{\mathbf{H}_{e}}(\theta_{2})\cdot K_{2}$ for $\theta_{2}>0$ as defined in Lemma 4.1. Now for any leaf $\mathcal{F}_{q}=\mathbf{H}_{e}(\mathbb{R})\cdot q$ ($q\in\mathbf{F}_{e}(\mathbb{Q})$) in the foliation $\mathcal{F}_{\mathbf{H}_{e}}$ such that
$\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})\neq\emptyset,$
we devide the region $\mathcal{F}_{q}\cap S$ into small cubes of side length
$\theta_{1}\cdot\exp(-\nu_{0}(a_{T_{k}}))/10$
where $\theta_{1}>0$ is the constant defined in Lemma 4.1. Then we collect those cubes which intersect $S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})$, and denote the corresponding collection by
$\mathcal{G}_{k,q,S}=\{R:R\cap S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})\neq\emptyset\}.$
Note that $\theta_{1}\cdot\exp(-\nu_{0}(a_{T_{k}}))/10$ is smaller than the minimum side length of the rectangle $B_{\mathbf{H}_{e}}(\theta_{1},T_{k})$, so by Lemma 4.1, we know that each cube in $\mathcal{G}_{k,q,S}$ is contained in
$S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})$
where $\tilde{K}_{2}=B_{\mathbf{H}_{e}}(\theta_{2})\cdot K_{2}$. Let
$\mathcal{H}_{k,S}=\bigcup_{\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})\neq\emptyset}\bigcup_{R\in\mathcal{G}_{k,q,S}}R.$
Then we have
$l_{k}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(S)$
$\sim m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}}))$
$\leq m_{\mathbf{H}_{e}}(\bigcup_{\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})\neq\emptyset}\bigcup_{R\in\mathcal{G}_{k,q,S}}R)$
$\leq m_{\mathbf{H}_{e}}(S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}}))$
$\sim l_{k}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(S)$
and
$m_{\mathbf{H}_{e}}(\mathcal{H}_{k,S})\sim l_{k}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(S).$
Note that each cube in $\mathcal{H}_{k,S}$ is contained in $S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})$. Let
$t_{k}=\frac{\ln l_{k}}{\beta_{0}(a_{-1})-(\tau+\epsilon)}.$
By the computations in §3, we can choose a sufficiently small number $\tilde{\delta}_{0}>0$ such that
$B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\times\exp(\mathfrak{a}_{K_{1},I_{0}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma\to B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\exp(\mathfrak{a}_{K_{1},I_{0}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma$
is a homeomorphism, and the subsets in the collection
$\mathcal{F}_{k}(S):=\left\{(a_{-t_{k}}\cdot B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\cdot a_{t_{k}})\cdot q:q\in\mathcal{H}_{k,S}\right\}$
are disjoint by this homeomorphism as
$a_{-t_{k}}\cdot B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\cdot a_{t_{k}}\subset B_{\mathbf{F}_{e}}(\tilde{\delta}_{0},T_{k}).$
We denote by
$\mathcal{P}_{k,S}=\bigcup_{E\in\mathcal{F}_{k}(S)}E.$
Now we can divide the subset $\mathcal{P}_{k,S}$ into cubes of side length
$\tilde{\delta}_{0}\cdot l_{k}^{\frac{\nu_{0}(a_{-1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}}$
which is smaller than $\theta_{1}\cdot\exp(-\nu_{0}(a_{T_{k}}))/10$ and the minimum side length of the rectangle
$a_{-t_{k}}\cdot B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\cdot a_{t_{k}}$
if $T_{k}>0$ and $l_{k}>0$ are sufficiently large. Note that $\Phi^{0}(\mathbf{F}_{e})$ may not be empty, and some of these cubes may be outside $S$, and here we collect only those cubes which are inside $S$. Thus, we obtain a family of disjoint cubes constructed from sets in $\mathcal{F}_{k}(S)$ inside $S$. We denote by $\mathcal{A}_{k}$ the collection of all these disjoint cubes constructed from $\mathcal{F}_{k}(S)$ inside $S$ where $S$ ranges over all elements in $\mathcal{A}_{k-1}$.
In this manner, we obtain a sequence $\left\{l_{k}\right\}_{k\in\mathbb{N}}$ of sufficiently large numbers $l_{k}$ with
$l_{k+1}\gg l_{k}\;(\forall k\in\mathbb{N})$
and a tree-like structure $\left\{\mathcal{A}_{k}\right\}_{k\in\mathbb{N}}$ of finite collections of cubes. Using the notation in Theorem 4.3, we have
$d_{k}(\mathcal{A})\sim l_{k}^{\frac{\nu_{0}(a_{-1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}}$
where $d_{k}(\mathcal{A})$ is the diameter of the family $\mathcal{A}_{k}$. Moreover, one can compute that
$\begin{cases}\Delta_{k}(\mathcal{A})\sim l_{k+1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})}{\beta_{0}(a_{1})}}\cdot\prod_{\alpha\in\Phi(\mathbf{F}_{e})}l_{k+1}^{\frac{\alpha(a_{-1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}},&\Phi^{0}(\mathbf{F}_{e})=\emptyset\\ \Delta_{k}(\mathcal{A})\sim l_{k+1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})}{\beta_{0}(a_{1})}}\cdot\prod_{\alpha\in\Phi^{1}(\mathbf{F}_{e})}l_{k+1}^{\frac{\alpha(a_{-1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}}\cdot\prod_{\alpha\in\Phi^{0}(\mathbf{F}_{e})}l_{k}^{\frac{\nu_{0}(a_{-1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}},&\Phi^{0}(\mathbf{F}_{e})\neq\emptyset.\end{cases}$
Now let $\mathbf{A}_{\infty}=\bigcap_{k\in\mathbb{N}}\mathbf{A}_{k}$. By applying Theorem 4.3, we can compute (assuming that $l_{k+1}$ is much larger than $l_{k}$ for any $k\in\mathbb{N}$) that
$\dim_{H}(\mathbf{A}_{\infty})\geq$ $\dim_{H}X-\limsup_{k\to\infty}\frac{\sum_{i=0}^{k}\log(\Delta_{i}(\mathcal{A}))}{\log(d_{k+1}(\mathcal{A}))}$
$=$ $\dim R_{u}(\overline{\mathbf{P}}_{0})-\sum_{\alpha\in\Phi^{1}(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}}$
$=$ $\dim R_{u}(\overline{\mathbf{P}}_{0})-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}}.$
Lemma 4.4.
We have $\mathbf{A}_{\infty}\subset S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
Proof.
By the construction, for any $p\in\mathbf{A}_{\infty}$, there exists a sequence of rational points
$q_{k}\in S_{K_{1},K_{2},K_{3},K_{4},j}(S_{k-1},A_{l_{k}},B_{l_{k}}),\quad S_{k-1}\in\mathcal{A}_{k-1}$
such that
$p\in(a_{-t_{k}}\cdot B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\cdot a_{t_{k}})\cdot q_{k}$
where $t_{k}=\frac{\ln l_{k}}{\beta_{0}(a_{-1})-(\tau+\epsilon)}$. Then
$a_{t_{k}}\cdot p\in B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\cdot(a_{t_{k}}\cdot q_{k}).$
Note that by definition, the denominator of the rational element $a_{t_{k}}\cdot q_{k}$ is equal to the co-volume of the lattice $a_{t_{k}}\cdot q_{k}\mathbb{Z}^{d}\cap V_{\beta_{0}}$ in $V_{\beta_{0}}$ and
$d(a_{t_{k}}\cdot q_{k})=e^{\beta_{0}(a_{t_{k}})\cdot\dim V_{\beta_{0}}}\cdot d(q_{k}).$
By Lemma 2.6, we deduce that
$\delta(\rho(a_{t_{k}}\cdot p)\mathbb{Z}^{d})$ $\sim\delta(\rho(a_{t_{k}}\cdot q_{k})\mathbb{Z}^{d})\lesssim d(a_{t_{k}}\cdot q_{k})^{\frac{1}{\dim V_{\beta_{0}}}}$
$\lesssim e^{\beta_{0}(a_{t_{k}})}\cdot l_{k}\leq e^{-(\tau+\epsilon)t_{k}}$
and $p\in S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. This completes the proof of the lemma. ∎
By Lemma 4.4 and the computation for $\dim_{H}\mathbf{A}_{\infty}$, we have
$\dim_{H}(S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))\geq\dim R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}}.$
By taking $\epsilon\to 0$, we obtain
$\dim_{H}(S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))\geq$ $\dim R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-\tau}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}}.$
Using the same argument as in [17, Section 10], we conclude that
$\dim_{H}(S_{\tau}^{c})\geq$ $\dim\mathbf{G}(\mathbb{R})-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-\tau}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}}.$
Note that by Lemma 4.2, we have $\nu_{0}(a_{1})>0$ and the argument here works for all $0\leq\tau<\beta_{0}(a_{-1})$. This completes the proof of Theorem 1.3 (provided that $\{a_{t}\}_{t\in\mathbb{R}}$ is mixing on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$). ∎
5. Proof of Theorem 1.3: $\{a_{t}\}_{t\in\mathbb{R}}$ in a proper normal $\mathbb{Q}$-subgroup of $\mathbf{G}$
In this section, we discuss the case when the action of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$ is not mixing, and explain how to modify the arguments in §3 and §4 and give a proof of Theorem 1.3.
We first discuss the ergodic properties of group actions on homogeneous spaces, and one can refer to e.g. [29, 35, 37] for details. Let $\mathbf{G}_{i}$ $(1\leq i\leq k)$ be the $\mathbb{Q}$-simple factors of $\mathbf{G}$. Then $\mathbf{G}$ is an almost direct product of $\mathbf{G}_{i}$ $(1\leq i\leq k)$. Without loss of generality, we may assume that $\{a_{t}\}_{t\in\mathbb{R}}$ projects nontrivially into $\mathbf{G}_{i}(\mathbb{R})^{0}$ ($1\leq i\leq s$) for some $s<k$. It is known that for each $1\leq i\leq k$, any arithmetic lattice $\Gamma_{i}$ inside $\mathbf{G}_{i}(\mathbb{Z})\cap\mathbf{G}_{i}(\mathbb{R})^{0}$ is irreducible in $\mathbf{G}_{i}(\mathbb{R})^{0}$ (and we fix such an arithmetic lattice $\Gamma_{i}$ for later use). Moreover, since $\{a_{t}\}_{t\in\mathbb{R}}$ projects nontrivially into $\mathbf{G}_{i}(\mathbb{R})^{0}$ $(1\leq i\leq s)$, we have $\mathbf{T}\cap\mathbf{G}_{i}\neq\{e\}$ $(1\leq i\leq s)$. So $\mathbf{G}_{i}$ is $\mathbf{Q}$-isotropic, and $\mathbf{G}_{i}(\mathbb{R})^{0}/\Gamma_{i}$ is not compact. By Godement compactness criterion (Cf. [7, Theorem 11.6]) and [8, 6.21], every simple factor of $\mathbf{G}_{i}(\mathbb{R})^{0}$ is not compact. Let $\{a_{t}^{i}\}_{t\in\mathbb{R}}$ be the projection of $\{a_{t}\}_{t\in\mathbb{R}}$ in $\mathbf{G}_{i}(\mathbb{R})^{0}$ $(1\leq i\leq s)$. Then $\{a_{t}^{i}\}_{t\in\mathbb{R}}\subset\mathbf{T}(\mathbb{R})\cap\mathbf{G}_{i}(\mathbb{R})$ is a non-compact subgroup in $\mathbf{G}_{i}(\mathbb{R})^{0}$. Using Moore’s ergodicity theorem [33] and Mautner phenomenon [30, 34], one can conclude that the action of $\{a_{t}^{i}\}_{t\in\mathbb{R}}$ on $\mathbf{G}_{i}(\mathbb{R})^{0}/\Gamma_{i}$ is mixing $(1\leq i\leq s)$. Consequently, the action of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\prod_{i=1}^{s}\mathbf{G}_{i}(\mathbb{R})^{0}/\prod_{i=1}^{s}\Gamma_{i}$ is mixing.
We denote by
$\tilde{\mathbf{G}}=\prod_{i=1}^{s}\mathbf{G}_{i}=\mathbf{G}_{1}\cdot\mathbf{G}_{2}\cdot\cdots\cdot\mathbf{G}_{s}.$
Note that in §2, §3 and §4, we don’t use any explicit expression of the lattice $\Gamma$ in $\mathbf{G}(\mathbb{R})^{0}$. So we may choose $\Gamma$ such that
$\Gamma\cap\tilde{\mathbf{G}}(\mathbb{R})^{0}=\prod_{i=1}^{s}\Gamma_{i}.$
In the follwing, we fix these lattices $\Gamma_{i}$’s and $\Gamma$ and denote by
$\tilde{\Gamma}=\Gamma\cap\tilde{\mathbf{G}}(\mathbb{R})^{0}.$
Therefore, the action of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\tilde{\mathbf{G}}(\mathbb{R})^{0}/\tilde{\Gamma}$ is mixing.
Now all the arguments in §2, §3 and §4 can be carried over to $\tilde{\mathbf{G}}$ and the homogeneous subspace $\tilde{\mathbf{G}}(\mathbb{R})^{0}/\tilde{\Gamma}$ almost verbatim. Indeed, denote by
$\tilde{\mathbf{P}}_{0}=\mathbf{P}_{0}\cap\tilde{\mathbf{G}}$
which is a minimal parabolic $\mathbb{Q}$-subgroup of $\tilde{\mathbf{G}}$, and let $\overline{\tilde{\mathbf{P}}}_{0}=\overline{\mathbf{P}}_{0}\cap\tilde{\mathbf{G}}$. Let $U$ be an open bounded subset in $R_{u}(\overline{\tilde{\mathbf{P}}}_{0})(\mathbb{R})$ and define
$\tilde{S}(U,A,B)=\{q\in U:q\textup{ rational and }A\leq d(q)\leq B\}.$
Let $\tilde{\mathbf{T}}=\mathbf{T}\cap\tilde{\mathbf{G}}$ which is a maximal $\mathbb{Q}$-split torus in $\tilde{\mathbf{G}}$, and
$\tilde{\mathbf{H}}_{e}=\mathbf{H}_{e}\cap\tilde{\mathbf{G}},\quad\tilde{\mathbf{F}}_{e}=\mathbf{F}_{e}\cap\tilde{\mathbf{G}}.$
In the following, we need results in the reduction theory about $\tilde{\mathbf{G}}(\mathbb{R})/\tilde{\Gamma}$. Let $\tilde{K}$ be a maximal compact subgroup in $\tilde{\mathbf{G}}(\mathbb{R})$. Let $\tilde{M}$ be the connected component of identity in the unique maximal $\mathbb{Q}$-anisotropic subgroup in $Z_{\tilde{\mathbf{G}}(\mathbb{R})}(\tilde{\mathbf{T}}(\mathbb{R}))$ (= the centralizer of $\tilde{\mathbf{T}}(\mathbb{R})$ in $\tilde{\mathbf{G}}(\mathbb{R})$). Denote by
$\tilde{T}_{\eta}=\{a\in\tilde{\mathbf{T}}(\mathbb{R}):\lambda(a)\leq\eta,\;\lambda\textup{ a simple root in }\tilde{\Delta}\}$
where $\tilde{\Delta}$ is the set of positive $\mathbb{Q}$-simple roots in $\tilde{\mathbf{G}}$. A Siegel set in $\tilde{\mathbf{G}}(\mathbb{R})$ is a subset of the form $\tilde{S}_{\eta,\Omega}=\tilde{K}\cdot\tilde{T}_{\eta}\cdot\tilde{\Omega}$ for some $\eta\in\mathbb{R}$ and a relatively compact open subset $\tilde{\Omega}$ containing identity in $\tilde{M}\cdot R_{u}(\tilde{\mathbf{P}}_{0})(\mathbb{R})$, and the group $\tilde{\mathbf{G}}(\mathbb{R})$ can be written as
$\tilde{\mathbf{G}}(\mathbb{R})=\tilde{S}_{\eta,\Omega}\cdot\tilde{\mathcal{K}}\cdot\tilde{\Gamma}$
for some Siegel set $\tilde{S}_{\eta,\Omega}$ and some finite subset $\tilde{\mathcal{K}}\subset\tilde{\mathbf{G}}(\mathbb{Q})$. Moreover, the finite set $\tilde{\mathcal{K}}$ satisfies the property that
$\tilde{\mathbf{G}}(\mathbb{Q})=\tilde{\mathbf{P}}_{0}(\mathbb{Q})\cdot\tilde{\mathcal{K}}\cdot\tilde{\Gamma}.$
Denote by $\tilde{\mathcal{K}}=\{\tilde{x}_{j}\}_{j\in\tilde{J}}\subset\tilde{\mathbf{G}}(\mathbb{Q})$ and we may assume that $e\in\tilde{\mathcal{K}}$.
Lemma 5.1.
We have
  1. (1)
    An element $g\in R_{u}(\overline{\tilde{\mathbf{P}}}_{0})(\mathbb{R})$ is rational if and only if $g\in\tilde{\mathbf{H}}_{e}(\mathbb{R})\cdot\tilde{\mathbf{F}}_{e}(\mathbb{Q})$.
  2. (2)
    An element $g\in R_{u}(\overline{\tilde{\mathbf{P}}}_{0})(\mathbb{R})$ is rational if and only if
    $g\in R_{u}(\overline{\tilde{\mathbf{P}}}_{0})\cap(\tilde{\mathbf{H}}_{1}(\mathbb{R})\cdot\tilde{\mathbf{P}}_{0}(\mathbb{R})\cdot\tilde{\mathcal{K}}\cdot\tilde{\Gamma}).$
Definition 5.2.
Let $j\in\tilde{J}$. A rational element $g$ in $R_{u}(\overline{\tilde{\mathbf{P}}}_{0})(\mathbb{R})$ is called $j$-rational if it can be written as
$g=h\cdot p\cdot\tilde{x}_{j}\cdot\gamma$
for some $h\in\tilde{\mathbf{H}}_{e}(\mathbb{R})$, $p\in\tilde{\mathbf{P}}_{0}(\mathbb{R})$ and $\gamma\in\tilde{\Gamma}$.
Let $\tilde{\mathfrak{a}}$ be the Lie algebra of $\tilde{\mathbf{T}}(\mathbb{R})$. For any $\tilde{x}_{j}\in\tilde{\mathcal{K}}$, any compact subset $K_{1}$ in $\ker(\beta_{0})\cap\tilde{\mathfrak{a}}$, any compact subset $K_{2}$ in $\tilde{\mathbf{H}}_{e}(\mathbb{R})$, any compact subset $K_{3}\subset\mathbf{M}_{a}(\mathbb{R})\cap\tilde{\mathbf{G}}(\mathbb{R})^{0}$ and any compact subset $K_{4}\subset R_{u}(\tilde{\mathbf{P}}_{0})(\mathbb{R})$, we define
$\tilde{S}_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)$
to be the set of all rational points $q$ in $U$ such that
  1. (1)
    $A\leq d(q)\leq B$ and $q$ is $j$-rational for $\tilde{x}_{j}\in\tilde{C}$;
  2. (2)
    $q=a\cdot h\cdot m\cdot u\cdot\tilde{x}_{j}\cdot\gamma$ for some $a\in\exp(\tilde{\mathfrak{a}})$, $\pi_{\ker(\beta_{0})}(a)\in K_{1}$, $h\in K_{2}$, $m\in K_{3}$, $u\in K_{4}$ and $\gamma\in\tilde{\Gamma}$.
Note that $\tilde{S}_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)\neq\emptyset$ implies that $\tilde{x}_{j}\in\tilde{\mathbf{G}}(\mathbb{R})^{0}$.
We fix a Haar measure $\mu_{\tilde{\mathbf{H}}_{e}}$ on $\tilde{\mathbf{H}}_{e}(\mathbb{R})$ and a Haar measure $\mu_{\tilde{\mathbf{F}}_{e}}$ on $\tilde{\mathbf{F}}_{e}(\mathbb{R})$. Then for any $q\in\tilde{\mathbf{F}}_{e}(\mathbb{Q})$, we define $m_{\tilde{\mathbf{H}}_{e}q}$ to be the locally finite measure defined on $R_{u}(\overline{\tilde{\mathbf{P}}}_{0})(\mathbb{R})$ which is supported on $\tilde{\mathbf{H}}_{e}(\mathbb{R})\cdot q$ and induced by $\mu_{\tilde{\mathbf{H}_{e}}}$ via the product map
$\tilde{\mathbf{H}}_{e}\times\{q\}\to\tilde{\mathbf{H}}_{e}\cdot q\subset R_{u}(\overline{\tilde{\mathbf{P}}}_{0}).$
We define
$m_{\tilde{\mathbf{H}}_{e}}:=\sum_{q\in\tilde{\mathbf{F}}_{e}(\mathbb{Q})}m_{\tilde{\mathbf{H}}_{e}q}.$
Then using the same argument as in Proposition 3.3, we have
Proposition 5.3.
Let $U$ be a small open bounded subset in $R_{u}(\overline{\tilde{\mathbf{P}}}_{0})$. Then for any $j\in\tilde{J}$ with $\tilde{x}_{j}\in\tilde{\mathbf{G}}(\mathbb{R})^{0}$ and any sufficiently large $l>0$, we have
$m_{\tilde{\mathbf{H}}_{e}}(\tilde{S}_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))\sim l^{-\sum_{\alpha\in\Phi(\tilde{\mathbf{F}}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\tilde{\mathbf{P}}}_{0})}(U)$
and
$m_{\tilde{\mathbf{H}}_{e}}(\tilde{S}_{K_{1},K_{2},K_{3},K_{4}}(U,A_{l},B_{l}))\sim l^{-\sum_{\alpha\in\Phi(\tilde{\mathbf{F}}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\tilde{\mathbf{P}}}_{0})}(U).$
Here the implicit constants depend only on the compact subsets $K_{i}$’s, $\tilde{\mathbf{G}}$ and $\tilde{\Gamma}$.
The analogues of Lemma 4.1 and Lemma 4.2 hold as well in $\tilde{\mathbf{G}}(\mathbb{R})^{0}/\tilde{\Gamma}$.
Lemma 5.4.
Fix $j\in\tilde{J}$ with $\tilde{x}_{j}\in\tilde{\mathbf{G}}(\mathbb{R})^{0}$. Let $U$ be an open bounded subset in $R_{u}(\overline{\tilde{\mathbf{P}}}_{0})(\mathbb{R})$ and
$\beta_{0}(a_{T})=-\ln l$
for some $T>0$ and $l>1$. Let ~$\tilde{\mathcal{F}}_{q}=\tilde{\mathbf{H}}_{e}(\mathbb{R})\cdot q$ be the leaf through $q\in\tilde{\mathbf{F}}_{e}(\mathbb{Q})$ such that
~$\tilde{\mathcal{F}}_{q}\cap\tilde{S}_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\neq\emptyset.$
Then there exist $\theta_{1}>0$ and $\theta_{2}>0$ such that for any $p\in\tilde{\mathcal{F}}_{q}\cap\tilde{S}_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$
$B_{\tilde{\mathbf{H}}_{e}}(\theta_{1},T)\cdot p\cap U\subset\tilde{\mathcal{F}}_{q}\cap\tilde{S}_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$
where $\tilde{K}_{2}=B_{\tilde{\mathbf{H}}_{1}}(\theta_{2})\cdot K_{2}$ and $B_{\tilde{\mathbf{H}}_{1}}(\theta_{1},T)=a_{-T}\cdot B_{\tilde{\mathbf{H}}_{e}}(\theta_{1})\cdot a_{T}$. Here the constants $\theta_{1}$ and $\theta_{2}$ depend only on $K_{i}$’s, $\tilde{\mathbf{G}}$ and $\tilde{\Gamma}$.
Lemma 5.5.
We have
$\max\{\alpha(a_{1}):\alpha\in\Phi(\tilde{\mathbf{F}}_{e})\}>0.$
Consequently, using the same argument as in §4, we can conclude that
$\dim_{H}(S_{\tau}^{c}\cap\tilde{\mathbf{G}}(\mathbb{R}))\geq\dim\tilde{\mathbf{G}}-\sum_{\alpha\in\Phi(\tilde{\mathbf{F}}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-\tau}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\tilde{\nu}_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}}$
where $\tilde{\nu}_{0}$ is the $\mathbb{Q}$-root in $R_{u}(\overline{\tilde{\mathbf{P}}}_{0})$ such that
$\tilde{\nu}_{0}(a_{1})=\max\{\alpha(a_{1}):\alpha\in\Phi(R_{u}(\overline{\tilde{\mathbf{P}}}_{0}))\}.$
Note that $\mathbf{G}_{i}(\mathbb{R})$ $(s+1\leq i\leq k)$ all commute with $\{a_{t}\}_{t\in\mathbb{R}}$, and an element $g\in S_{\tau}^{c}$ if and only if $h\cdot g\in S_{\tau}^{c}$ for any $h\in\mathbf{G}_{i}(\mathbb{R})$ $(s+1\leq i\leq k)$. Therefore, we have
$\dim_{H}(S_{\tau}^{c})=$ $\dim_{H}(S_{\tau}^{c}\cap\tilde{\mathbf{G}}(\mathbb{R}))+\sum_{i=s+1}^{k}\dim\mathbf{G}_{i}(\mathbb{R})$
$\geq$ $\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-\tau}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}}$
where $\nu_{0}(a_{1}):=\max\{\alpha(a_{1}):\alpha\in\Phi(\mathbf{F}_{e})\}$. Here we use the fact that $\nu_{0}(a_{1})=\tilde{\nu}_{0}(a_{1})$ and
$\sum_{\alpha\in\Phi(\tilde{\mathbf{F}}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})=\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1}).$
This completes the proof of Theorem 1.3.
Remark 5.6.
One can see from the arguments in §4 and §5 that we actually prove that for any open bounded subset $U$ in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$
$\dim_{H}(S_{\gamma}^{c}\cap U)\geq\dim R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-\tau}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}}.$
Note that $\mathbf{P}_{0}\cdot R_{u}(\overline{\mathbf{P}}_{0})$ is a Zariski open dense subset in $\mathbf{G}$, and if $x\in S_{\gamma}^{c}\cap U$, then for any $g\in\mathbf{P}_{0}(\mathbb{R})$, $g\cdot x$ is also an element in $S_{\gamma}^{c}$. This implies that for any open bounded subset $\tilde{U}\subset\mathbf{G}(\mathbb{R})$
$\dim_{H}(S_{\gamma}^{c}\cap\tilde{U})\geq\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-\tau}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}}.$
6. An upper bound for the Hausdorff dimension of $S_{\gamma}^{c}$
In this section, we prove Theorem 1.4. We will compute an upper bound of $\dim_{H}S_{\gamma}^{c}$ by constructing open covers of the subset $S_{\gamma}^{c}$. Recall that
$\mathcal{C}_{\bar{I}}:=\bigcup_{i\in\bar{I}}w_{i}\mathbf{F}_{w_{i}}(\mathbb{R}).$
Proposition 6.1.
Suppose that
$\delta(\rho(g)\cdot\mathbb{Z}^{d})\leq r$
for some $g\in\mathbf{G}(\mathbb{R})$ and $r>0$. Then there exist $\tilde{g}\in\mathcal{C}_{\bar{I}}$ and a discrete primitive subgroup
$\Lambda_{g}\subset\rho(g)\cdot\mathbb{Z}^{d}$
such that
$\operatorname{Vol}(\Lambda_{g})\leq C\cdot r^{\dim V_{\beta_{0}}}\textup{ and }\rho(\tilde{g})\cdot\Lambda_{g}\subset V_{\beta_{0}}\textup{ is Zariski densein }V_{\beta_{0}}$
where the constant $C>0$ depends only on $\mathbf{G}$, $\Gamma$ and $\rho$.
Proof.
By the reduction theory of arithmetic subgroups, $\mathbf{G}(\mathbb{R})$ can be written as
$\mathbf{G}(\mathbb{R})=S_{\eta,\Omega}\cdot\mathcal{K}\cdot\Gamma$
for some Siegel set $S_{\eta,\Omega}=K\cdot T_{\eta}\cdot\Omega$ and some finite subset $\mathcal{K}\subset\mathbf{G}(\mathbb{Q})$ where
$T_{\eta}=\{a\in T(\mathbb{R}):\lambda(a)\leq\eta,\;\lambda\textup{ a simple root in }\Delta\}$
and the finite set $\mathcal{K}$ satisfies the property that
$\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot\mathcal{K}\cdot\Gamma.$
Then there exist $k\in K$, $a\in T_{\eta}$, $u\in\Omega$, $\gamma\in\Gamma$ and $x\in\mathcal{K}\subset\mathbf{G}(\mathbb{Q})$ such that
$g=k\cdot a\cdot u\cdot x\cdot\gamma.$
Note that by the definition of $T_{\eta}$, the subset
$\{aua^{-1}:a\in T_{\eta}\textup{ and }u\in\Omega\}$
is compact. So we can write
$g=\tilde{k}\cdot a\cdot x\cdot\gamma$
for some $\tilde{k}$ in a fixed compact subset in $\mathbf{G}(\mathbb{R})$.
Now since $\delta(\rho(g)\cdot\mathbb{Z}^{d})\leq r$, there exists $y\in\mathbb{Z}^{d}\setminus\{0\}$ such that $\rho(\tilde{k}\cdot a\cdot x)\cdot y$ is a shortest vector in $\rho(g)\cdot\mathbb{Z}^{d}$ and
$\|\rho(a\cdot x)\cdot y\|\sim\|\rho(\tilde{k}\cdot a\cdot x)\cdot y\|\leq r.$
We write
$\rho(x)\cdot y=\sum_{\text{weight }\beta}y_{\beta}$
according to the decomposition of the complex vector space $V=\oplus_{\beta}V_{\beta}$ in $\rho$ into weight spaces $V_{\beta}$ relative to $\mathbf{T}$, where $y_{\beta}\in V_{\beta}$. Let $\tilde{\beta}$ be a weight among the weights $\beta$’s in the summation such that $y_{\tilde{\beta}}\neq 0$. Then by the structure of the representation $\rho$, we compute that
(1) $\|\rho(a\cdot x)\cdot y\|=\|e^{\tilde{\beta}(a)}y_{\tilde{\beta}}+\cdots\|\lesssim r.$
Note that $y\in\mathbb{Z}^{d}$ and $x\in\mathcal{K}$, and there is a positive lower bound for $\|y_{\tilde{\beta}}\|$ depending only on $\rho$ and $\mathcal{K}$. On the other hand, there exists a unique discrete primitive subgroup $\tilde{\Lambda}\subset\mathbb{Z}^{d}$ such that $\rho(x)\cdot\tilde{\Lambda}$ is a discrete and Zariski-dense subgroup in $V_{\beta_{0}}$, and for any $w\in\tilde{\Lambda}\setminus\{0\}$ we have
(2) $\|\rho(\tilde{k}\cdot a\cdot x)\cdot w\|\sim\|\rho(a)\cdot\rho(x)\cdot w\|=e^{\beta_{0}(a)}\|\rho(x)\cdot w\|.$
Comparing equations (1) and (2) and using the fact that $x\in\mathcal{K}\subset\mathbf{G}(\mathbb{Q})$, $a\in T_{\eta}$ and the relation between $\tilde{\beta}$ and $\beta_{0}$ according to the structure of the representation $\rho$, one can deduce that
$e^{\beta_{0}(a)}\lesssim_{\eta}e^{\tilde{\beta}(a)}\lesssim r$
and for any $w\in\tilde{\Lambda}\setminus\{0\}$
$\|\rho(\tilde{k}\cdot a\cdot x)w\|\lesssim r\cdot\|\rho(x)w\|.$
Let $\Lambda_{g}=\rho(g)(\rho(\gamma^{-1})\tilde{\Lambda})$. Then we have
$\Lambda_{g}\subset\rho(g)\cdot\mathbb{Z}^{d},\quad\operatorname{Vol}(\Lambda_{g})\lesssim r^{\operatorname{rank}\Lambda_{g}}=r^{\dim V_{\beta_{0}}}\textup{ and }\rho(\tilde{k}^{-1})\cdot\Lambda_{g}\subset V_{\beta_{0}}.$
By the Bruhat decomposition, we know that
$\tilde{k}^{-1}\in\bigcup_{i\in\bar{I}}\mathbf{P}_{\beta_{0}}(\mathbb{R})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{R}).$
By the fact that $\mathbf{P}_{\beta_{0}}$ stabilizes $V_{\beta_{0}}$, we conclude that there exists $\tilde{g}\in\mathcal{C}_{\bar{I}}$ such that
$\tilde{k}^{-1}\in\mathbf{P}_{\beta_{0}}(\mathbb{R})\cdot\tilde{g}\textup{ and }\rho(\tilde{g})\cdot\Lambda_{g}\subset V_{\beta_{0}}.$
This completes the proof of the proposition. ∎
Let $\epsilon_{0}>0$ which we will determine later. Let $g$ be an element in $S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Then by definition, there exists a sequence $\{t_{k}\}\subset\mathbb{N}$ such that
$\delta(\rho(a_{t_{k}}g)\cdot\mathbb{Z}^{d})\leq\epsilon_{0}\cdot e^{-\tau t_{k}}.$
By Proposition 6.1, for each $k\in\mathbb{N}$, there exist $\tilde{g}_{k}\in\mathcal{C}_{\bar{I}}$ and a discrete subgroup $\Lambda_{k}\subset\rho(a_{t_{k}}g)\cdot\mathbb{Z}^{d}$ such that
$\operatorname{Vol}(\Lambda_{k})\lesssim(\epsilon_{0}e^{-\tau t_{k}})^{\dim V_{\beta_{0}}}\textup{ and }\rho(\tilde{g}_{k})\cdot\Lambda_{k}\textup{ is Zariski dense in }V_{\beta_{0}}.$
Let $\tilde{g}_{k}=w_{i,k}\cdot f_{i,k}\in\mathcal{C}_{\bar{I}}$ where $w_{i,k}\in\{w_{i}\}_{i\in\bar{I}}$ and $f_{i,k}\in\mathbf{F}_{w_{i,k}}(\mathbb{R})$. Then it implies that
$\tilde{q}_{k}:=w_{i,k}\cdot f_{i,k}\cdot a_{t_{k}}\cdot g=w_{i,k}\cdot a_{t_{k}}\cdot(a_{-t_{k}}f_{i,k}a_{t_{k}})\cdot g$
$q_{k}:=(w_{i,k}a_{-t_{k}}w_{i,k}^{-1})\cdot\tilde{q}_{k}=w_{i,k}(a_{-t_{k}}f_{i,k}a_{t_{k}})g$
are rational elements in $\mathbf{G}(\mathbb{R})$. In the $\dim V_{\beta_{0}}$-exterior product space of $V$, we can compute that
$d(q_{k})=e^{\beta_{0}(w_{i,k}a_{-t_{k}}w_{i,k}^{-1})\cdot\dim V_{\beta_{0}}}\cdot d(\tilde{q}_{k})$
$d(\tilde{q}_{k})\cdot e_{V_{\beta_{0}}}=\rho_{\beta_{0}}(\tilde{g}_{k})\cdot\left(\bigwedge^{\dim V_{\beta_{0}}}\Lambda_{k}\right),\quad\rho_{\beta_{0}}(\tilde{g}_{k}^{-1})\cdot e_{V_{\beta_{0}}}=\left(\bigwedge^{\dim V_{\beta_{0}}}\Lambda_{k}\right)/d(\tilde{q}_{k})$
$\|\rho_{\beta_{0}}(\tilde{g}_{k}^{-1})\cdot e_{V_{\beta_{0}}}\|=$ $\operatorname{Vol}(\Lambda_{k})/d(\tilde{q}_{k})$
$\lesssim$ $(\epsilon_{0}\cdot e^{-\tau t_{k}})^{\dim V_{\beta_{0}}}\cdot e^{\beta_{0}(w_{i,k}a_{-t_{k}}w_{i,k}^{-1})\cdot\dim V_{\beta_{0}}}/d(q_{k})$
where $e_{V_{\beta_{0}}}=\bigwedge^{\dim V_{\beta_{0}}}V_{\beta_{0}}$ is the unit vector which represents the vector space $V_{\beta_{0}}$ in the $\dim V_{\beta_{0}}$-exterior product space of $V$, and $\bigwedge^{\dim V_{\beta_{0}}}\Lambda_{k}$ is the vector constructed from a $\mathbb{Z}$-basis of $\Lambda_{k}$ and represents the lattice $\Lambda_{k}$. The implicit constants here depend only on $\mathbf{G}$, $\Gamma$ and $\rho$.
Definition 6.2.
Fix an arbitrary $i\in\bar{I}$. For any $R>0$, define
$E_{w_{i}}(R):=\{f\in\mathbf{F}_{w_{i}}(\mathbb{R}):\|\rho_{\beta_{0}}(w_{i}fw_{i}^{-1})\cdot e_{V_{\beta_{0}}}\|\leq R\}.$
We also define the following morphism
$\Psi_{w_{i}}:\mathbf{F}_{w_{i}}(\mathbb{R})\to\bigwedge_{j=1}^{\dim V_{\beta_{0}}}V,\quad\Psi_{w_{i}}(x)=\rho_{\beta_{0}}(w_{i}xw_{i}^{-1})\cdot e_{V_{\beta_{0}}}.$
Note that $w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}\subset R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$. Hence $\Psi_{w_{i}}$ is an isomorphism onto its image.
From the discussion above, we know that
$\|\rho_{\beta_{0}}(w_{i,k}f_{i,k}^{-1}w_{i,k}^{-1})\cdot e_{V_{\beta_{0}}}\|\lesssim$ $\|\rho_{\beta_{0}}(\tilde{g}_{k}^{-1})\cdot e_{V_{\beta_{0}}}\|$
$\lesssim$ $(\epsilon_{0}\cdot e^{-\tau t_{k}})^{\dim V_{\beta_{0}}}\cdot e^{\beta_{0}(w_{i,k}a_{-t_{k}}w_{i,k}^{-1})\cdot\dim V_{\beta_{0}}}/d(q_{k})$
and by definition of $E_{w_{i,k}}(R)$, we have
$g=$ $(a_{-t_{k}}f_{i,k}^{-1}a_{t_{k}})\cdot w_{i,k}^{-1}\cdot q_{k}$
$\in$ $\left(a_{-t_{k}}\cdot E_{w_{i,k}}\left(C_{1}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i,k}a_{-t_{k}}w_{i,k}^{-1})})^{\dim V_{\beta_{0}}}/d(q_{k})\right)\cdot a_{t_{k}}\right)\cdot w_{i,k}^{-1}\cdot q_{k}$
for some constant $C_{1}>0$ depending only on $\mathbf{G}$, $\Gamma$ and $\rho$. Moreover, by the structure of the representation $\rho$, we know that
$\|\rho_{\beta_{0}}(w_{i,k}f_{i,k}^{-1}w_{i,k}^{-1})\cdot e_{V_{\beta_{0}}}\|=\|\Psi_{w_{i,k}}(f_{i,k}^{-1})\|\gtrsim\|e_{V_{\beta_{0}}}\|=1$
so
$d(q_{k})\leq C_{2}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i,k}a_{-t_{k}}w_{i,k}^{-1})})^{\dim V_{\beta_{0}}}$
for some constant $C_{2}>0$ depending only on $\mathbf{G}$, $\Gamma$ and $\rho$. Therefore, $q_{k}$ is a rational element in $\mathbf{G}(\mathbb{R})$ with
$d(q_{k})\leq C_{2}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i,k}a_{-t_{k}}w_{i,k}^{-1})})^{\dim V_{\beta_{0}}}$
and $w_{i,k}^{-1}\cdot q_{k}\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. By passing to a subsequence, we may assume that $w_{i,k}$ is a fixed element in $\{w_{i}\}_{i\in\bar{I}}$ independent of $k$. In the following, we take $C_{0}=\max\{C_{1},C_{2}\}$.
Definition 6.3.
For any $i\in\bar{I}$, define $\mathcal{E}_{i}(\tau)$ to be the subset of elements $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ for which there exist a divergent sequence $\{t_{k}\}\subset\mathbb{N}$ and a sequence of rational elements $\{q_{k}\}\subset\mathbf{G}(\mathbb{R})$ such that
$g\in\left(a_{-t_{k}}E_{w_{i}}(C_{0}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i}a_{-t_{k}}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/d(q_{k}))a_{t_{k}}\right)w_{i}^{-1}q_{k}$
$d(q_{k})\leq C_{0}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i}a_{-t_{k}}w_{i}^{-1})})^{\dim V_{\beta_{0}}}$
and $w_{i}^{-1}q_{k}\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
Lemma 6.4.
Let $i\in\bar{I}$. Let $q$ be a rational element in $\mathbf{G}(\mathbb{R})$ with $w_{i}^{-1}\cdot q\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Then there exists a constant $\theta_{i}>0$ depending only on $i$ such that $d(q)\geq\theta_{i}$.
Proof.
By definition, we know that $R_{u}(\overline{\mathbf{P}}_{0})=\mathbf{H}_{w_{i}}\cdot\mathbf{F}_{w_{i}}.$ By Corollary 2.3, let
$w_{i}^{-1}\cdot q=u\cdot v$
for some $u\in\mathbf{H}_{w_{i}}(\mathbb{R})$ and $v\in\mathbf{F}_{w_{i}}(\mathbb{Q})$. Let $\tilde{\Omega}_{\mathbf{F}_{w_{i}}}$ be a bounded fundamental domain of $\mathbf{F}_{w_{i}}(\mathbb{R})/(\Gamma\cap\mathbf{F}_{w_{i}}(\mathbb{R}))$ in $\mathbf{F}_{w_{i}}(\mathbb{R})$. We may write
$v=\tilde{v}\cdot\gamma$
for some $\tilde{v}\in\tilde{\Omega}_{\mathbf{F}_{w_{i}}}$ and $\gamma\in\Gamma\cap\mathbf{F}_{w_{i}}(\mathbb{R})$. Then
$q\cdot\mathbb{Z}^{d}=(w_{i}\cdot u\cdot v)\cdot\mathbb{Z}^{d}=(w_{i}uw_{i}^{-1})\cdot(w_{i}\tilde{v}\cdot\mathbb{Z}^{d}).$
By the fact that $w_{i}\mathbf{H}_{w_{i}}w_{i}^{-1}$ fixes every element in $V_{\beta_{0}}$ and $\tilde{v}$ is in the bounded subset $\tilde{\Omega}_{\mathbf{F}_{w_{i}}}$, we can conclude that the co-volume of $(w_{i}\tilde{v}\cdot\mathbb{Z}^{d})\cap V_{\beta_{0}}$ has a lower bound depending only on $w_{i}$ and $\tilde{\Omega}_{\mathbf{F}_{w_{i}}}$, and so does $q\cdot\mathbb{Z}^{d}$. This completes the proof of the lemma. ∎
Now we choose $\epsilon_{0}>0$ so that
$\min_{i\in\bar{I}}\{\theta_{i}\}>C_{0}\cdot\epsilon_{0}^{\dim V_{\beta_{0}}}.$
Lemma 6.5.
If $\mathcal{E}_{i}(\tau)\neq\emptyset$ for some $i\in\bar{I}$, then $0\leq\tau<\beta_{0}(w_{i}a_{-1}w_{i}^{-1})$.
Proof.
Let $g\in\mathcal{E}_{i}(\tau)$. By definition, there exist a divergent sequence $\{t_{k}\}\subset\mathbb{N}$ and a sequence of rational elements $\{q_{k}\}\subset\mathbf{G}(\mathbb{R})$ such that
$g\in\left(a_{-t_{k}}E_{w_{i}}(C_{0}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i}a_{-t_{k}}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/d(q_{k}))a_{t_{k}}\right)w_{i}^{-1}q_{k}$
$d(q_{k})\leq C_{0}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i}a_{-t_{k}}w_{i}^{-1})})^{\dim V_{\beta_{0}}}$
and $w_{i}^{-1}q_{k}\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. By Lemma 6.4, we have
$\theta_{i}\leq C_{0}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i}a_{-t_{k}}w_{i}^{-1})})^{\dim V_{\beta_{0}}}.$
If $\tau\geq\beta_{0}(w_{i}a_{-1}w_{i}^{-1})$, then by taking $t_{k}\to\infty$, we obtain that
$\theta_{i}\leq C_{0}\cdot(\epsilon_{0})^{\dim V_{\beta_{0}}}$
which contradicts the choice of $\epsilon_{0}$. This completes the proof of the lemma. ∎
From the discussion above, we obtain the following
Proposition 6.6.
We have
$S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\subseteq\bigcup_{i\in\bar{I}}\mathcal{E}_{i}(\tau).$
In the following, for each $i\in\bar{I}$, we will consider the subset $S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau)$ $(\mathcal{E}_{i}(\tau)\neq\emptyset)$ and compute an upper bound for the Hausdorff dimension of $S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau)$. These upper bounds will give an upper bound for $\dim_{H}(S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ by Proposition 6.6. For notational convenience, we may set $\epsilon_{0}=1$ as it does not affect the computations in the rest of this section.
Now fix $i\in\bar{I}$. To compute the Hausdorff dimension of $S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau)$, we first need to estimate the volume of the subset $E_{w_{i}}(R)$ in ${\mathbf{F}}_{w_{i}}(\mathbb{R})$, which can be written as
$\operatorname{Vol}(E_{w_{i}}(R))=\mu_{\mathbf{F}_{w_{i}}}(\Psi_{w_{i}}^{-1}(B_{R}))$
where $B_{R}$ is the ball of radius $R$ centered at 00 in $\bigwedge_{i=1}^{\dim V_{\beta_{0}}}V$ and $\mu_{\mathbf{F}_{w_{i}}}$ is the Haar measure on ${\mathbf{F}}_{w_{i}}(\mathbb{R})$. We need the following result about the asymptotic volume estimates of algebraic varieties.
Theorem 6.7 ([3, Corollary 16.3]).
[…]
$\mu(B_{R})\sim cR^{a}(\log R)^{b}\;(\textup{as }R\to\infty)$
for some $a\in\mathbb{Q}_{\geq 0}$, $b\in\mathbb{Z}_{\geq 0}$ and $c>0$.
We apply Theorem 6.7 to the morphism $\Psi_{w_{i}}$ where the closed orbit $\mathcal{O}=\Psi_{w_{i}}(\mathbf{F}_{w_{i}}(\mathbb{R}))$ and the measuer $\mu$ is the push-forward $\Psi_{w_{i}}^{*}(\mu_{\mathbf{F}_{w_{i}}})$ of the Haar measure on $\mathbf{F}_{w_{i}}(\mathbb{R})$, and obtain the following
Corollary 6.8.
There exist constants $a_{w_{i}}\in\mathbb{Q}_{\geq 0}$, $b_{w_{i}}\in\mathbb{Z}_{\geq 0}$ and $c_{w_{i}}>0$ such that
$\operatorname{Vol}(E_{w_{i}}(R))=\mu_{\mathbf{F}_{w_{i}}}(\Psi_{w_{i}}^{-1}(B_{R}))\sim c_{w_{i}}R^{a_{w_{i}}}(\log R)^{b_{w_{i}}}\;(\textup{as }R\to\infty).$
To compute the Hausdorff dimension of $S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau)$, we also need an upper bound for the asymptotic number of rational elements in
$(w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1})(\mathbb{Q})\subset R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q}).$
Consider the morphism
$\tilde{\Psi}_{w_{i}}:w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}(\mathbb{R})\to\bigwedge_{i=1}^{\dim V_{\beta_{0}}}V,\quad\tilde{\Psi}_{w_{i}}(x)=\rho_{\beta_{0}}(x)\cdot e_{V_{\beta_{0}}}.$
Let $g\in w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}(\mathbb{Q})\subset R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q})$ be a rational element in a bounded set $\tilde{U}$ in $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})$. By definition, $d(g)$ is the co-volume of $\rho(g)\mathbb{Z}^{d}\cap V_{\beta_{0}}$ in $V_{\beta_{0}}(\mathbb{R})$. In the $\dim V_{\beta_{0}}$-exterior product space of $V$, this implies that there exists a constant $C>0$ depending only on $\tilde{U}$, such that the length of the primitive integral vector in the line spanned by $\rho(g^{-1})\cdot e_{V_{\beta_{0}}}$ is less than $C\cdot d(g)$. So the number of rational points $g$ in $\tilde{U}$ whose denominators are less than $l>0$ is less than or equal to the total number of primitive integral points in $\bigwedge_{i=1}^{\dim V_{\beta_{0}}}V$ whose lengths are less than $C\cdot l$. This leads us to the results about the Manin’s conjecture [2] (for relevant results about the Manin’s conjecture, one may refer to [11, 16, 18, 19, 32, 36, 38, 40] and the references therein.) In particular, there exists a constant $A_{w_{i}}>0$ such that for any $\epsilon>0$
($\ast$) #|$\#\left|\{g\in\tilde{U}\cap w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}(\mathbb{Q}):d(g)\leq l\}\right|\ll_{\epsilon,\tilde{U}}l^{A_{w_{i}}+\epsilon}$
where the implicit constant depends only on $\epsilon$, $\tilde{U}$, $\mathbf{G}$ and $\rho$. In the following, we will fix the constants $a_{w_{i}},b_{w_{i}},c_{w_{i}}$ and $A_{w_{i}}>0$.
Lemma 6.9.
Let $i\in\bar{I}$ such that $\beta_{0}(w_{i}a_{-1}w_{i}^{-1})>0.$ Then
$\max\{\alpha(a_{1}):\alpha\in\Phi(\mathbf{F}_{w_{i}})\}>0.$
Proof.
The proof is similar to those of Lemma 4.2 and Lemma 5.5. Suppose on the contrary that for all $\alpha\in\Phi(\mathbf{F}_{w_{i}})$, $\alpha(a_{1})=0$. Let $\Omega_{\mathbf{F}_{w_{i}}}$ and $\Omega_{\mathbf{H}_{w_{i}}}$ be small open neighborhoods of identity in $\mathbf{F}_{w_{i}}(\mathbb{R})$ and $\mathbf{H}_{w_{i}}(\mathbb{R})$ respectively, and
$\Omega_{R_{u}(\overline{\mathbf{P}}_{0})}:=\Omega_{\mathbf{F}_{w_{i}}}\cdot\Omega_{\mathbf{H}_{w_{i}}}.$
Note that $w_{i}\mathbf{H}_{w_{i}}(\mathbb{R})w_{i}^{-1}$ stabilizes every element in $V_{\beta_{0}}$. Now for any $p=f\cdot h\in\Omega_{R_{u}(\overline{\mathbf{P}}_{0})}$ with $f\in\Omega_{\mathbf{F}_{w_{i}}}$ and $h\in\Omega_{\mathbf{H}_{w_{i}}}$, we have
$a_{t}\cdot p\cdot w_{i}^{-1}=(a_{t}fa_{-t})(a_{t}ha_{-t})a_{t}w_{i}^{-1}=fw_{i}^{-1}\cdot(w_{i}a_{t}ha_{-t}w_{i}^{-1})\cdot w_{i}a_{t}w_{i}^{-1}.$
The element $w_{i}a_{t}w_{i}^{-1}$ is rational and
$\delta(\rho(w_{i}a_{t}w_{i}^{-1})\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}})=e^{\beta_{0}(w_{i}a_{t}w_{i}^{-1})}.$
This implies that
$\delta(\rho(a_{t}\cdot p\cdot w_{i}^{-1})\cdot\mathbb{Z}^{d})\leq\kappa\cdot e^{-\beta_{0}(w_{i}a_{-1}w_{i}^{-1})t}$
for any $t>0$ where the constant $\kappa$ depends only on $\Omega_{\mathbf{F}_{w_{i}}}$ and $w_{i}$. Let $\Omega_{\mathbf{P}_{0}}$ be a small neighborhood of identity in $\mathbf{P}_{0}$. Then one can deduce from the inequality above that for any point $p\Gamma\in\Omega_{\mathbf{P}_{0}}\cdot\Omega_{R_{u}(\overline{\mathbf{P}}_{0})}\cdot w_{i}^{-1}\Gamma$, the orbit $a_{t}\cdot p\Gamma$ diverges in $\mathbf{G}(\mathbb{R})/\Gamma$. Set $\Gamma_{w_{i}}=\Gamma\cap w_{i}^{-1}\Gamma w_{i}$. Then $\Gamma_{w_{i}}$ is commensurable with $\Gamma$ and $w_{i}^{-1}\Gamma w_{i}$, and we can conclude that for any $p\Gamma_{w_{i}}\in\Omega_{\mathbf{P}_{0}}\cdot\Omega_{R_{u}(\overline{\mathbf{P}}_{0})}\cdot\Gamma_{w_{i}}$, the orbit $a_{t}\cdot p\Gamma_{w_{i}}$ diverges in $\mathbf{G}(\mathbb{R})^{0}/\Gamma_{w_{i}}$. Similiar to Lemma 4.2 and Lemma 5.5, this contradicts the ergodic property of the flow $\{a_{t}\}_{t\in\mathbb{R}}$ on the homogeneous subspace $\tilde{\mathbf{G}}(\mathbb{R})^{0}/\Gamma_{w_{i}}\cap\tilde{\mathbf{G}}(\mathbb{R})^{0}$, where $\tilde{\mathbf{G}}$ is the product of some $\mathbb{Q}$-simple factors $\mathbf{G}_{i}$ $(1\leq i\leq s)$ of $\mathbf{G}$ such that $\{a_{t}\}_{t\in\mathbb{R}}$ projects nontrivially into $\mathbf{G}_{i}(\mathbb{R})$ $(1\leq i\leq s)$. This completes the proof of the lemma. ∎
Now we compute an upper bound for the Hausdorff dimension of the subset $S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau)$. In the computation, we project the subset $S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau)$ into the quotient space $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ by the natural projection map
$\pi_{R_{u}(\overline{\mathbf{P}}_{0})}:R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\to R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
and compute the Hausdorff dimension of $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}(S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau))$. Since $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}$ is a local isometry, by the countable stability of Hausdorff dimension, we then obtain the upper bounds for $\dim_{H}(\pi_{R_{u}(\overline{\mathbf{P}}_{0})}^{-1}(\pi_{R_{u}(\overline{\mathbf{P}}_{0})}(S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau))))$ and $\dim_{H}(S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau))$.
Let $\tilde{\Omega}$ be a bounded fundamental domain of $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. We choose a bounded subset $\Omega_{\mathbf{H}_{w_{i}}}$ in $\mathbf{H}_{w_{i}}(\mathbb{R})$ and a bounded subset $\Omega_{\mathbf{F}_{w_{i}}}$ in $\mathbf{F}_{w_{i}}(\mathbb{R})$ such that
$\tilde{\Omega}\subset\Omega_{\mathbf{H}_{w_{i}}}\cdot\Omega_{\mathbf{F}_{w_{i}}}.$
Now suppose that $q\in\mathbf{G}(\mathbb{R})$ is a rational element with $w_{i}^{-1}\cdot q\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Let $\tilde{q}\in\mathbf{G}(\mathbb{R})$ such that $w_{i}^{-1}\tilde{q}$ is a representative of $w_{i}^{-1}q$ in $\tilde{\Omega}$ with
$w_{i}^{-1}\tilde{q}=w_{i}^{-1}q\gamma$
for some $\gamma\in\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Then one can check that $\tilde{q}$ is still a rational element in $\mathbf{G}(\mathbb{R})$ and $d(\tilde{q})=d(q)$ since $\Gamma$ preserves the lattice $\mathbb{Z}^{d}$. So by Corollary 2.3, we still have
$w_{i}^{-1}\tilde{q}\in\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{F}_{w_{i}}(\mathbb{Q}).$
Since $w_{i}^{-1}\tilde{q}$ is an element in the bounded fundamental domain $\tilde{\Omega}$, it implies that
$w_{i}^{-1}\tilde{q}\in\Omega_{\mathbf{H}_{w_{i}}}\cdot(\mathbf{F}_{w_{i}}(\mathbb{Q})\cap\Omega_{\mathbf{F}_{w_{i}}}).$
We then obtain the following
Proposition 6.10.
The set $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}(S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau))$ consists of points $g(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ for which there exist a divergent sequence $\{t_{k}\}\subset\mathbb{N}$ and a sequence of rational elements $\{q_{k}\}\subset\mathbf{G}(\mathbb{R})$ such that
$g(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))\in\left(a_{-t_{k}}E_{w_{i}}(C_{0}(e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i}a_{-t_{k}}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/d(q_{k}))a_{t_{k}}\right)w_{i}^{-1}q_{k}(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
$d(q_{k})\leq C_{0}(e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}$
and $w_{i}^{-1}q_{k}\in\tilde{\Omega}$.
Lemma 6.11.
Let $\delta>0$ and $B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta)$ be the small open ball of radius $\delta$ centered at identity in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Then there exists a constant $C_{\delta}>0$ depending only on $\delta$ such that for any $R>0$
$B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta)\cdot E_{w_{i}}(R)\subset E_{w_{i}}(C_{\delta}\cdot R)\cdot\mathbf{H}_{w_{i}}(\mathbb{R}).$
Proof.
Let $x\in B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta)$ and $y\in E_{w_{i}}(R)$. Then $x\cdot y\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. We write
$x\cdot y=u\cdot v$
where $u\in\mathbf{F}_{w_{i}}(\mathbb{R})$ and $v\in\mathbf{H}_{w_{i}}(\mathbb{R})$. Then we have
$\|\Psi_{w_{i}}(u)\|=$ $\|\rho_{\beta_{0}}(w_{i}uw_{i}^{-1})\cdot e_{V_{\beta_{0}}}\|=\|\rho(w_{i}uvw_{i}^{-1})\cdot e_{V_{\beta_{0}}}\|$
$=$ $\|\rho(w_{i}xyw_{i}^{-1})\cdot e_{V_{\beta_{0}}}\|\leq C_{\delta}\cdot R$
for some constant $C_{\delta}>0$ depending only on $\delta>0$. This completes the proof of the lemma. ∎
Note that by definition, for any rational element $g\in\mathbf{G}(\mathbb{R})$, every element $x$ in the $\Gamma$-coset $g\Gamma$ of $g$ is also rational, and $d(x)=d(g)$. We will use this fact in the following lemma.
Lemma 6.12.
Let $i\in\bar{I}$ and $L>0$. Then the set of points $g(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ with $w_{i}\cdot g$ rational and $d(w_{i}\cdot g)=L$ is equal to
$\{(x\cdot y)(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})):x\in\Omega_{\mathbf{H}_{w_{i}}},y\in\Omega_{\mathbf{F}_{w_{i}}}\cap\mathbf{F}_{w_{i}}(\mathbb{Q}),d(w_{i}\cdot y)=L\}.$
Proof.
Let $g(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))\in R_{u}(\overline{\mathbf{P}}_{0})/(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0}))$ with $w_{i}\cdot g$ rational and $d(w_{i}\cdot g)=L$. We write
$g=(u\cdot v)\cdot\gamma$
for some $u\in\Omega_{\mathbf{H}_{w_{i}}}$, $v\in\Omega_{\mathbf{F}_{w_{i}}}$ and $\gamma\in\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})$. Then we have
$w_{i}\cdot g=(w_{i}\cdot u\cdot w_{i}^{-1})\cdot w_{i}\cdot v\cdot\gamma$
and $w_{i}\cdot v$ is rational with
$d(w_{i}\cdot g)=d(w_{i}\cdot v)=L.$
By Corollary 2.3, we know that $v\in\mathbf{F}_{w_{i}}(\mathbb{Q})$. This completes the proof of the lemma. ∎
Remark 6.13.
For any $L>0$, we define
$\mathcal{S}(L):=\{(x\cdot y)(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})):x\in\Omega_{\mathbf{H}_{w_{i}}},y\in\Omega_{\mathbf{F}_{w_{i}}}\cap\mathbf{F}_{w_{i}}(\mathbb{Q}),d(w_{i}\cdot y)=L\}$
$\mathcal{F}(L):=\{y\in\Omega_{\mathbf{F}_{w_{i}}}\cap\mathbf{F}_{w_{i}}(\mathbb{Q}):d(w_{i}\cdot y)=L\}.$
Now we construct open covers of $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}(S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau))$ with arbitrarily small diameters. Here we use the right-invariant metric $d_{R_{u}(\overline{\mathbf{P}}_{0})}$ on $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ which induces a metric $d_{R_{u}(\overline{\mathbf{P}}_{0})/\Gamma}$ on the quotient space $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$. Note that by Corollary 6.8, for any $\epsilon>0$ and $R>0$, we have
$\operatorname{Vol}(E_{w_{i}}(R))\lesssim_{\epsilon}R^{a_{w_{i}}+\epsilon}$
where the implicit constant depends only on $\epsilon$ and $i\in\bar{I}$. Let $\nu_{0}$ be the $\mathbb{Q}$-root in $R_{u}(\overline{\mathbf{P}}_{0})$ such that
$\nu_{0}(a_{1})=\max\{\alpha(a_{1}):\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{0}))\}.$
By Lemma 6.9, we have $\nu_{0}(a_{1})>0$.
Fix $\delta>0$. Let $L>0$ and $t\in\mathbb{N}$ such that
$L\leq C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}.$
Let $q$ be a rational element in $\mathbf{G}(\mathbb{R})$ such that $d(q)=L$ and $w_{i}^{-1}q\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}).$ The following subset
$\left(a_{-t}\cdot E_{w_{i}}\left(C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/d(q)\right)\cdot a_{t}\right)\cdot w_{i}^{-1}q\cdot(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
can be covered by disjoint boxes of diameter at most $\delta\cdot\exp(\nu_{0}(a_{-t}))$, and by Lemma 6.11 and Lemma 6.12, we know that these boxes are contained in
$B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta\cdot\exp(\nu_{0}(a_{-t})))\left(a_{-t}\cdot E_{w_{i}}\left(C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/d(q)\right)\cdot a_{t}\right)\cdot w_{i}^{-1}q\cdot(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
$\subset$ $\left(a_{-t}\cdot B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta)\cdot E_{w_{i}}\left(C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/d(q)\right)\cdot a_{t}\right)\cdot w_{i}^{-1}q\cdot(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
$\subset$ $\left(a_{-t}\cdot E_{w_{i}}\left(C_{\delta}\cdot C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/L\right)\cdot a_{t}\right)\cdot\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathcal{S}(L).$
Since every point $g(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ in $\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathcal{S}(L)$ satisfies the property that $w_{i}\cdot g$ is rational and $d(w_{i}\cdot g)=L$, it follows again from Lemma 6.12 that
$\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathcal{S}(L)=\mathcal{S}(L).$
We conclude that the subset
$\left(a_{-t}E_{w_{i}}(C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/d(q))a_{t}\right)\cdot w_{i}^{-1}q\cdot(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
can be covered by disjoint boxes with diameter at most $\delta\cdot\exp(\nu_{0}(a_{-t}))$ and these boxes are contained in
$a_{-t}E_{w_{i}}\left(C_{\delta}\cdot C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/L\right)a_{t}\cdot\mathcal{S}(L).$
We collect all these boxes constructed above in a family which is denoted by $\mathcal{P}_{t,L}(q)$, and define
$\mathcal{P}_{t,L}=\bigcup_{d(q)=L\textup{ and }w_{i}^{-1}q\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})}\mathcal{P}_{t,L}(q).$
Now since every box in $\mathcal{P}_{t,L}$ is contained in the same bounded set
$a_{-t}E_{w_{i}}(C_{\delta}\cdot C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/L)a_{t}\cdot\mathcal{S}(L)$
we can choose a maximal finite sub-collection $\mathcal{Q}_{t,L}$ of disjoint boxes in $\mathcal{P}_{t,L}$. Then by the maximality of $\mathcal{Q}_{t,L}$ we have
$B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta\cdot\exp(\nu_{0}(a_{-t})))^{2}\cdot\mathcal{Q}_{t,L}\supset\bigcup_{S\in\mathcal{P}_{t,L}}S.$
Meanwhile, the number of boxes in $\mathcal{Q}_{t,L}$ is at most
$\frac{\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(a_{-t}E_{w_{i}}(C_{\delta}\cdot C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/L)a_{t})\cdot\mathcal{S}(L))}{\exp(\dim R_{u}(\overline{\mathbf{P}}_{0})\cdot\nu_{0}(a_{-t}))}.$
Now we define $\mathcal{G}_{t}$ to be the collection of boxes in
$B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta\cdot\exp(\nu_{0}(a_{-t})))^{2}\cdot\mathcal{Q}_{t,L}$
where
$\theta_{i}\leq L\leq C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}.$
Note that $\operatorname{diam}\mathcal{G}_{t}\sim e^{\nu_{0}(a_{-t})}$ and $\nu_{0}(a_{-t})\neq 0$ by Lemma 6.9. Define
$\mathcal{G}^{k}=\bigcup_{t\geq k}\mathcal{G}_{t}.$
Then by definition and Proposition 6.10, the subset $\mathcal{G}^{k}$ is a cover of $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}(S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau))$ for any $k\in\mathbb{N}$ and $\operatorname{diam}\mathcal{G}^{k}\sim e^{\nu_{0}(a_{-k})}$.
Now according to the construction of the open covers $\mathcal{G}^{k}$ $(k\in\mathbb{N})$ above, we consider the following series with respect to the parameter $s$
$\sum_{B\in\mathcal{G}^{k}}\operatorname{diam}(B)^{s}\leq\sum_{t\in\mathbb{N}}\sum_{B\in\mathcal{G}_{t}}\operatorname{diam}(B)^{s}$
$\lesssim$ $\sum_{t\in\mathbb{N}}\frac{\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(a_{-t}E_{w_{i}}(C_{\delta}\cdot C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/L)a_{t})\cdot\mathcal{S}(L))}{\exp(\dim R_{u}(\overline{\mathbf{P}}_{0})\cdot\nu_{0}(a_{-t}))}\cdot e^{s\nu_{0}(a_{-t})}.$
By Corollary 6.8, for any $\epsilon>0$, we have
$\sum_{B\in\mathcal{G}^{k}}\operatorname{diam}(B)^{s}$
$\lesssim$ $\sum_{t\in\mathbb{N}}\sum_{\theta_{i}\leq L\leq C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}}(C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/L)^{a_{w_{i}}+\epsilon}\cdot|\mathcal{F}(L)|$
$e^{\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\alpha(a_{-t})}e^{-\dim R_{u}(\overline{\mathbf{P}}_{0})\cdot\nu_{0}(a_{-t})}\cdot e^{s\nu_{0}(a_{-t})}.$
Fix $t\in\mathbb{N}$. Note that $\tau<\beta_{0}(w_{i}a_{-1}w_{i}^{-1})$. Choose $t_{0}\in\mathbb{Z}$ such that
$C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot t_{0}})^{\dim V_{\beta_{0}}}\leq\theta_{i}.$
Then by equation ($\ast$6) before Lemma 6.9, we have
$\sum_{\theta_{i}\leq L\leq C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}}|\mathcal{F}(L)|/L^{a_{w_{i}}+\epsilon}$
$\lesssim$ $\sum_{t_{0}\leq l\leq t-1}\sum_{C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot l})^{\dim V_{\beta_{0}}}\leq L\leq C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot(l+1)})^{\dim V_{\beta_{0}}}}|\mathcal{F}(L)|/L^{a_{w_{i}}+\epsilon}$
$\lesssim$ $\sum_{t_{0}\leq l\leq t-1}\frac{1}{(C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot l})^{\dim V_{\beta_{0}}})^{a_{w_{i}}+\epsilon}}\sum_{L\leq C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot(l+1)})^{\dim V_{\beta_{0}}}}|\mathcal{F}(L)|$
$\lesssim$ $\sum_{t_{0}\leq l\leq t-1}\frac{1}{(C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot l})^{\dim V_{\beta_{0}}})^{a_{w_{i}}+\epsilon}}(C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot(l+1)})^{\dim V_{\beta_{0}}})^{A_{w_{i}}+\epsilon}$
$\sim$ $\sum_{t_{0}\leq l\leq t-1}(C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot l})^{\dim V_{\beta_{0}}})^{A_{w_{i}}-a_{w_{i}}}.$
Combining all these equations, we obtain that
$\sum_{B\in\mathcal{G}^{k}}\operatorname{diam}(B)^{s}$
$\lesssim$ $\sum_{t\in\mathbb{N}}(C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)t})^{\dim V_{\beta_{0}}})^{a_{w_{i}}+\epsilon}\cdot e^{\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\alpha(a_{-t})}e^{-\dim R_{u}(\overline{\mathbf{P}}_{0})\cdot\nu_{0}(a_{-t})}\cdot e^{s\nu_{0}(a_{-t})}$
$\cdot\sum_{t_{0}\leq l\leq t-1}(C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot l})^{\dim V_{\beta_{0}}})^{A_{w_{i}}-a_{w_{i}}}.$
By computing the series above in the cases $A_{w_{i}}<a_{w_{i}}$, $A_{w_{i}}=a_{w_{i}}$ and $A_{w_{i}}>a_{w_{i}}$, one can conclude that the series above converges if
$s>\dim R_{u}(\overline{\mathbf{P}}_{0})-\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot(\max\{A_{w_{i}},a_{w_{i}}\}+\epsilon)\cdot\dim V_{\beta_{0}}.$
This implies that for any $\epsilon>0$
$\dim_{H}(S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau))\leq\dim R_{u}(\overline{\mathbf{P}}_{0})-\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot(\max\{A_{w_{i}},a_{w_{i}}\}+\epsilon)\cdot\dim V_{\beta_{0}}.$
By taking $\epsilon\to 0$ and Proposition 6.6, we conclude that
$\dim_{H}(S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
$\leq\max_{i\in\bar{I}}\left\{\dim R_{u}(\overline{\mathbf{P}}_{0})-\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{A_{w_{i}},a_{w_{i}}\}\cdot\dim V_{\beta_{0}}\right\}$
and
$\dim_{H}S_{\tau}^{c}\leq\max_{i\in\bar{I}}\left\{\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{A_{w_{i}},a_{w_{i}}\}\cdot\dim V_{\beta_{0}}\right\}.$
This completes the proof of Theorem 1.4.
7. Proofs of Theorems 1.61.7 and 1.8
In this section, we prove Theorems 1.61.7 and 1.8.
Proof of Theorem 1.6.
Let $\mathbf{T}$ be the full diagonal group in $\mathbf{G}=\operatorname{\mathbf{SL}}_{n}$. Without loss of generality, we may write
$\{a_{t}\}_{t\in\mathbb{R}}=\left\{\operatorname{diag}(e^{b_{1}t},e^{b_{2}t},\cdots,e^{b_{n}t}):t\in\mathbb{R}\right\}$
where $b_{1}\geq b_{2}\geq\cdots\geq b_{n}$ and $b_{1}+b_{2}+\cdots+b_{n}=0$. We write
$\mathbf{T}(\mathbb{R})^{0}=\{\operatorname{diag}(e^{t_{1}},e^{t_{2}},\cdots,e^{t_{n}}):t_{1}+t_{2}+\cdots+t_{n}=0\}$
where $\mathbf{T}(\mathbb{R})^{0}$ is the connected component of identity in $\mathbf{T}(\mathbb{R})$. Let $\mathbf{P}_{0}$ be the lower triangular subgroup in $\mathbf{G}$ and $R_{u}(\overline{\mathbf{P}}_{0})$ is the upper triangular unipotent subgroup in $\mathbf{G}$. Then all the $\mathbb{Q}$-roots in $\mathbf{G}$ with respect to $\mathbf{T}$ are
$\alpha_{i,j}(a)=t_{i}-t_{j}\quad(1\leq i\neq j\leq n)$
where $a=\operatorname{diag}(e^{t_{1}},e^{t_{2}},\cdots,e^{t_{n}})$. Let $\{e_{1},e_{2},...,e_{n}\}$ be the standard basis in $V$ where
$e_{1}=\left(\begin{array}[]{c}1\\ 0\\ 0\\ \vdots\\ 0\end{array}\right)\quad e_{2}=\left(\begin{array}[]{c}0\\ 1\\ 0\\ \vdots\\ 0\end{array}\right)\quad\cdots\cdots\quad e_{n}=\left(\begin{array}[]{c}0\\ 0\\ \vdots\\ 0\\ 1\end{array}\right).$
We choose $V_{\lambda_{0}}=V_{\beta_{0}}=\mathbb{C}\cdot e_{n}$ to be the highest weight space and
$V=\mathbb{C}\cdot e_{1}\oplus\mathbb{C}\cdot e_{2}\oplus\cdots\oplus\mathbb{C}\cdot e_{n}$
is the weight space decomposition of $V$. The stabilizer $\mathbf{P}_{\beta_{0}}$ of $V_{\beta_{0}}$ and $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ are
$\mathbf{P}_{\beta_{0}}=\left(\begin{array}[]{cccc}*&*&\cdots&0\\ &*&\cdots&0\\ \vdots&\vdots&\cdots&\vdots\\ &*&\cdots&*\end{array}\right)\textup{ and }R_{u}(\overline{\mathbf{P}}_{\beta_{0}})=\left(\begin{array}[]{cc}I_{n-1}&*\\ 0&1\end{array}\right).$
Suppose that $x\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q})$ is a rational element in $\mathbf{G}(\mathbb{R})$. Let
$x=\left(\begin{array}[]{cccc}1&0&\cdots&\frac{p_{1}}{q}\\ 0&1&\cdots&\frac{p_{2}}{q}\\ \vdots&\vdots&\cdots&\vdots\\ 0&0&\cdots&1\end{array}\right)$
where $p_{i}\in\mathbb{Z},q\in\mathbb{N}$ and $(p_{1},p_{2},\dots,p_{n-1},q)=1$. Then $x\cdot\mathbb{Z}^{n}\cap V_{\beta_{0}}\neq\{0\}$ implies that there exists a primitive integral point $(k_{1},k_{2},\dots,k_{n-1},l)^{T}\in\mathbb{Z}^{n}\setminus\{0\}$ such that
$\left(\begin{array}[]{cccc}1&0&\cdots&\frac{p_{1}}{q}\\ 0&1&\cdots&\frac{p_{2}}{q}\\ \vdots&\vdots&\cdots&\vdots\\ 0&0&\cdots&1\end{array}\right)\cdot\left(\begin{array}[]{c}k_{1}\\ k_{2}\\ \vdots\\ l\end{array}\right)=\left(\begin{array}[]{c}0\\ 0\\ \vdots\\ \end{array}\right).$
Then we have
$k_{i}+\frac{p_{i}}{q}\cdot l=0,\quad 1\leq i\leq n-1.$
Hence $|l|=q$ and the denominator $d(x)$ of $x$ is equal to $q$. One can then deduce from the formula $d(x)=q$ that for any bounded open subset $\tilde{U}$ in $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})$ and any sufficiently large $l>0$, the number of rational elements in $\tilde{U}$ whose denominators are less than $l$ is bounded by $\mu_{R_{u}(\overline{\mathbf{P}}_{\beta_{0}})}(\tilde{U})\cdot l^{n}.$ Therefore, we may choose $A_{e}=n$ in Equation ($\ast$6) in §6. One can also compute that $a_{e}=n-1$.
Now consider the upper bound of the Hausdorff dimension of $S_{\tau}^{c}$. In the computation, we actually use the Bruhat decomposition as
$\mathbf{G}(\mathbb{R})=\bigcup_{i\in\bar{I}}\mathbf{P}_{\beta_{0}}(\mathbb{R})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{R}).$
In our case, we may choose the index set $\bar{I}$ as small as possible (as the double cosets in the original decomposition with respect to Weyl group ${}_{\mathbb{Q}}W$ overlap). Note that ${}_{\mathbb{Q}}W$ is isomorphic to the symmetric group of permutations of $\{1,2,\dots,n\}$ and many elements of ${}_{\mathbb{Q}}W$ are inside $\mathbf{P}_{\beta_{0}}(\mathbb{R})$. From this observation, one can deduce that if $w_{i}=(i,n)$ $(1\leq i\leq n)$ is the permutation of $i$ and $n$, then
$\mathbf{G}(\mathbb{R})=\bigcup_{i=1}^{n}\mathbf{P}_{\beta_{0}}(\mathbb{R})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{R}).$
Then by definition, one may compute that $A_{w_{i}}\leq A_{e}$, $a_{w_{i}}\leq a_{e}$ for all $1\leq i\leq n$, and
$\mathbf{F}_{w_{i}}=\left(\begin{array}[]{cccccc}1&0&\cdots&*&\cdots&0\\ 0&1&\cdots&*&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&1&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&0&\cdots&1\end{array}\right)$
where the only non-zero entries are in the $i$-th column.
Now we compute the upper bound of $\dim_{H}S_{\tau}^{c}$. Fix $w_{i}$ $(1\leq i\leq n)$. Note that the roots in $\mathbf{F}_{w_{i}}$ are $\alpha_{1,i},\alpha_{2,i},\dots,\alpha_{i-1,i}$. According to Theorem 1.4 and the discussion in §6, we may assume that $\beta_{0}(w_{i}a_{-1}w_{i}^{-1})>0$, i.e. $b_{i}<0$; otherwise there is nothing to compute. Then $\nu_{0}=\alpha_{1,n}$ and by the fact that $b_{1}\geq b_{2}\geq\cdots\geq b_{n}$ and $b_{1}+b_{2}+\cdots+b_{n}=0$ we have
$\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{A_{w_{i}},a_{w_{i}}\}\cdot\dim V_{\beta_{0}}$
$=$ $\dim\mathbf{G}-\frac{(b_{1}-b_{i})+(b_{2}-b_{i})+\cdots+(b_{i-1}-b_{i})}{b_{1}-b_{n}}+\frac{-b_{i}-\tau}{b_{1}-b_{n}}\cdot\max\{A_{w_{i}},a_{w_{i}}\}$
$\leq$ $\dim\mathbf{G}-\frac{(b_{1}-b_{i})+(b_{2}-b_{i})+\cdots+(b_{i-1}-b_{i})}{b_{1}-b_{n}}+n\cdot\frac{-b_{i}-\tau}{b_{1}-b_{n}}$
$=$ $\dim\mathbf{G}-\frac{(b_{1}-b_{i})+(b_{2}-b_{i})+\cdots+(b_{i-1}-b_{i})+nb_{i}}{b_{1}-b_{n}}-\frac{n\tau}{b_{1}-b_{n}}$
$=$ $\dim\mathbf{G}-\frac{-b_{i+1}\cdots-b_{n}+(n-i)b_{i}}{b_{1}-b_{n}}-\frac{n\tau}{b_{1}-b_{n}}$
$=$ $\dim\mathbf{G}-\frac{(b_{i}-b_{i+1})+\cdots(b_{i}-b_{n})}{b_{1}-b_{n}}-\frac{n\tau}{b_{1}-b_{n}}$
$\leq$ $\dim\mathbf{G}-\frac{n\tau}{b_{1}-b_{n}}.$
Therefore, by Theorem 1.4, we have
$\dim_{H}S_{\tau}^{c}\leq\dim\mathbf{G}-\frac{n\cdot\tau}{\nu_{0}(a_{1})}.$
On the other hand, for the lower bound of $\dim_{H}S_{\tau}^{c}$, by Theorem 1.3, we have
$\dim_{H}S_{\tau}^{c}\geq$ $\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})\cdot\frac{\tau}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}$
$=$ $\dim\mathbf{G}-((b_{1}-b_{n})+(b_{2}-b_{n})+\cdots+(b_{n-1}-b_{n}))\cdot\frac{\tau}{(-b_{n})(b_{1}-b_{n})}$
$=$ $\dim\mathbf{G}-\frac{n\tau}{b_{1}-b_{n}}=\dim\mathbf{G}-\frac{n\cdot\tau}{\nu_{0}(a_{1})}.$
We conclude that
$\dim_{H}S_{\tau}^{c}=\dim\mathbf{G}-\frac{n\cdot\tau}{\nu_{0}(a_{1})}.$
This completes the proof of Theorem 1.6. ∎
Proof of Theorem 1.7.
Let $\mathbf{T}$ be the full diagonal group in $\mathbf{G}=\operatorname{\mathbf{SL}}_{n}$ and we may write
$\{a_{t}\}_{t\in\mathbb{R}}=\left\{\operatorname{diag}(e^{b_{1}t},e^{b_{2}t},\cdots,e^{b_{n}t}):t\in\mathbb{R}\right\}$
where $b_{1}\geq b_{2}\geq\cdots\geq b_{n}$ and $b_{1}+b_{2}+\cdots+b_{n}=0$. We write
$\mathbf{T}(\mathbb{R})^{0}=\{\operatorname{diag}(e^{t_{1}},e^{t_{2}},\cdots,e^{t_{n}}):t_{1}+t_{2}+\cdots+t_{n}=0\}$
where $\mathbf{T}(\mathbb{R})^{0}$ is the connected component of identity in $\mathbf{T}(\mathbb{R})$. Let $\mathbf{P}_{0}$ be the lower triangular subgroup in $\mathbf{G}$ and $R_{u}(\overline{\mathbf{P}}_{0})$ is the upper triangular unipotent subgroup in $\mathbf{G}$. Let $V_{\beta_{0}}$ be the highest weight space in $\mathfrak{sl}_{n}$
$V_{\beta_{0}}=\left(\begin{array}[]{cccc}0&0&\cdots&0\\ \vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&0\\ &0&\cdots&0\end{array}\right).$
Note that here $\nu_{0}=-\beta_{0}$ is the highest root in $\mathfrak{sl}_{n}$ and $\beta_{0}(a_{t})=(b_{n}-b_{1})t<0$. The stabilizer $\mathbf{P}_{\beta_{0}}$ of $V_{\beta_{0}}$ and $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ are
$\mathbf{P}_{\beta_{0}}=\left(\begin{array}[]{ccccc}*&0&\cdots&0&0\\ &*&\cdots&*&0\\ \vdots&\vdots&\cdots&\vdots&\vdots\\ &*&\cdots&*&0\\ &*&\cdots&*&*\end{array}\right)\textup{ and }R_{u}(\overline{\mathbf{P}}_{\beta_{0}})=\left(\begin{array}[]{ccccc}1&*&\cdots&*&*\\ 0&1&\cdots&0&*\\ \vdots&\vdots&\cdots&\vdots&\vdots\\ 0&0&\cdots&1&*\\ 0&0&\cdots&0&1\end{array}\right).$
Let $x\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q})$ be a rational element in $\mathbf{G}(\mathbb{R})$. We write
$x=\left(\begin{array}[]{ccccc}1&\frac{a_{1}}{b}&\cdots&\frac{a_{n-2}}{b}&0\\ 0&1&\cdots&0&0\\ \vdots&\vdots&\cdots&\vdots&\vdots\\ 0&0&\cdots&1&0\\ 0&0&\cdots&0&1\end{array}\right)\cdot\left(\begin{array}[]{ccccc}1&0&\cdots&0&\frac{p_{1}}{q}\\ 0&1&\cdots&0&\frac{p_{2}}{q}\\ \vdots&\vdots&\cdots&\vdots&\vdots\\ 0&0&\cdots&1&\frac{p_{n-1}}{q}\\ 0&0&\cdots&0&1\end{array}\right)$
where $(a_{1},a_{2},\dots,a_{n-2},b)=1$ and $(p_{1},p_{2},\dots,p_{n-1},q)=1$. Then by definition,
$x\cdot\mathfrak{sl}_{n}(\mathbb{Z})\cdot x^{-1}\cap V_{\beta_{0}}\neq\{0\}.$
It implies that there exists $l\in\mathbb{R}$ such that
$x^{-1}\cdot\left(\begin{array}[]{cccc}0&0&\cdots&0\\ \vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&0\\ l&0&\cdots&0\end{array}\right)\cdot x\in\mathfrak{sl}_{n}(\mathbb{Z}).$
One can compute that
$\left(\begin{array}[]{ccccc}-\frac{p_{1}}{q}\cdot l&-\frac{p_{1}}{q}\cdot l\cdot\frac{a_{1}}{b}&\cdots&-\frac{p_{1}}{q}\cdot l\cdot\frac{a_{n-2}}{b}&-\frac{p_{1}}{q}\cdot l\cdot\frac{bp_{1}+a_{1}p_{2}+\cdots+a_{n-2}p_{n-1}}{bq}\\ \vdots&\vdots&\cdots&\vdots&\vdots\\ -\frac{p_{n-1}}{q}\cdot l&-\frac{p_{n-1}}{q}\cdot l\cdot\frac{a_{1}}{b}&\cdots&-\frac{p_{n-1}}{q}\cdot l\cdot\frac{a_{n-2}}{b}&-\frac{p_{n-1}}{q}\cdot l\cdot\frac{bp_{1}+a_{1}p_{2}+\cdots+a_{n-2}p_{n-1}}{bq}\\ l&l\cdot\frac{a_{1}}{b}&\cdots&l\cdot\frac{a_{n-2}}{b}&l\cdot\frac{bp_{1}+a_{1}p_{2}+\cdots+a_{n-2}p_{n-1}}{bq}\end{array}\right)\in\mathfrak{sl}_{n}(\mathbb{Z})$
and consequently
$\gcd(p_{1},...,p_{n-1})\cdot\frac{l}{q},\;\gcd(p_{1},...,p_{n-1})\cdot\frac{l}{q}\cdot\frac{a_{i}}{b}(1\leq i\leq n-2),\;\gcd(p_{1},...,p_{n-1})\cdot\frac{l}{q}\cdot\frac{d}{bq}\in\mathbb{Z}$
$l,\;l\cdot\frac{a_{1}}{b},\dots,l\cdot\frac{a_{n-2}}{b},\;l\cdot\frac{d}{bq}\in\mathbb{Z}$
where $d=bp_{1}+a_{1}p_{2}+\cdots+a_{n-2}p_{n-1}$. Then one can deduce from the formulas above that
$l=s\cdot\frac{bq^{2}}{\gcd(d,q)}\;(s\in\mathbb{Z}).$
Therefore, the denominator $d(x)$ of $x$ is equal to
$d(x)=\frac{bq^{2}}{\gcd(d,q).}$
It follows from the same argument as in [17, Lemma 8.6] that
|$\left|\{x\in\tilde{U}:d(x)\leq l\}\right|\ll_{\epsilon,\tilde{U}}l^{n-1+\epsilon}$
for any bounded open subset $\tilde{U}$ in $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})$, and hence $A_{e}=n-1$.
Now we compute the upper bound of $\dim_{H}S_{\tau}^{c}$ using Theorem 1.4. We choose elements $\{w_{i}\}_{i\in\bar{I}}$ in the Weyl group ${}_{\mathbb{Q}}W$ such that the Bruhat decomposition stated in Theorem 1.4 holds
$\mathbf{G}(\mathbb{Q})=\bigcup_{i\in\bar{I}}\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{Q}).$
Note that many Weyl elements in ${}_{\mathbb{Q}}W$ are inside $\mathbf{P}_{\beta_{0}}$. Here we choose $\{w_{i}\}_{i\in\bar{I}}$ as follows. We identify ${}_{\mathbb{Q}}W$ with the symmetric group of permutations on $\{1,2,...,n\}$. Then there are three types of elements in $\{w_{i}\}_{i\in\bar{I}}$:
Type I:
$(1,2,\dots,n-1,n)$, $(n,2,\dots,n-1,1)$
Type II:
$(1,2\dots,j-1,n,j+1,\dots,n-1,j)$, $(j,2,\dots,j-1,n,j+1,\dots,n-1,1)$, $(j,2,\dots,j-1,1,j+1,\dots,n-1,n)$, $(n,2,\dots,j-1,1,j+1,\dots,n-1,j)$ ($12\leq j\leq n-$ 1)
Type III:
$(i,2,\dots i-1,1,i+1,\dots,j-1,n,j+1,\dots,n-1,j)$, $(j,2,\dots i-1,1,i+1,\dots,j-1,n,j+1,\dots n-1,i)$ ($12\leq i<j\leq n-$ 1).
The index set $\bar{I}$ has cardinality $n(n-1)$, and one may check that $\{w_{i}\}_{i\in\bar{I}}$ meets our requirement. Note that $A_{w_{i}}\leq A_{e}$ for any $i\in\bar{I}$.
For a Weyl element $w$ of type I in $\{w_{i}\}_{i\in\bar{I}}$, if $w=(1,2,\dots,n)$ is the identity element, then the number $a_{w}$ in Corollary 6.8 comes from the morphism $\Psi_{w}$ defined by
$\left(\begin{array}[]{ccccc}1&x_{1}&\cdots&x_{n-2}&z\\ 0&1&\cdots&0&y_{1}\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&1&y_{n-2}\\ 0&0&\cdots&0&1\end{array}\right)\cdot\left(\begin{array}[]{ccccc}0&0&\cdots&0&0\\ 0&0&\cdots&0&0\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&0&0\\ 1&0&\cdots&0&0\end{array}\right)\cdot\left(\begin{array}[]{ccccc}1&x_{1}&\cdots&x_{n-2}&z\\ 0&1&\cdots&0&y_{1}\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&1&y_{n-2}\\ 0&0&\cdots&0&1\end{array}\right)^{-1}$
$=$ $\left(\begin{array}[]{ccccc}z&-x_{1}\cdot z&\cdots&-x_{n-2}\cdot z&(x_{1}y_{1}+\cdots+x_{n-2}y_{n-2}-z)z\\ y_{1}&-x_{1}\cdot y_{1}&\cdots&-x_{n-2}\cdot y_{1}&(x_{1}y_{1}+\cdots+x_{n-2}y_{n-2}-z)y_{1}\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ y_{n-2}&-x_{1}\cdot y_{n-2}&\cdots&-x_{n-2}\cdot y_{n-2}&(x_{1}y_{1}+\cdots+x_{n-2}y_{n-2}-z)y_{n-2}\\ 1&-x_{1}&\cdots&-x_{n-2}&x_{1}y_{1}+\cdots+x_{n-2}y_{n-2}-z\end{array}\right)$
where $x_{1},x_{2},\dots,x_{n-2},y_{1},y_{2},\dots,y_{n-2},z$ are the variables of $\Psi_{w}$. One can estimate $\operatorname{Vol}(E_{w}(R))$ and compute that $a_{w}\leq n-1$. Then according to Theorem 1.4, for $w$, we have
$\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(wa_{-1}w^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{A_{w},a_{w}\}\cdot\dim V_{\beta_{0}}$
$=$ $\dim\mathbf{G}-(n-1)+\frac{(b_{1}-b_{n})-\tau}{b_{1}-b_{n}}\cdot(n-1)=\dim\mathbf{G}-\frac{(n-1)\tau}{\nu_{0}(a_{1})}.$
Note that for $w=(n,2,3,\cdots,n-1,1)$, $\beta_{0}(wa_{-1}w^{-1})\leq 0$ and there is nothing to compute in this case.
Now let us check a Weyl element $w$ of type II in $\{w_{i}\}_{i\in\bar{I}}$. If $w=(j,2,\dots,j-1,n,j+1,\dots,n-1,1)$ or $w=(n,2,\dots,j-1,1,j+1,\dots,n-1,j)$ ($12\leq j\leq n-$ 1), then $\beta_{0}(w\cdot a_{-1}\cdot w^{-1})\leq 0$ and there is nothing to compute. If $w=(1,2,\dots,j-1,n,j+1,\dots,n-1,j)$ $(2\leq j\leq n-1)$, one can compute that
$\mathbf{F}_{w}=\left(\begin{array}[]{ccccccc}1&*&*&\cdots&*&\cdots&*\\ 0&1&0&\cdots&*&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&0&\cdots&1&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&0&\cdots&0&\cdots&1\end{array}\right).$
One can then deduce that $a_{w}\leq n-1$, which is similar to the type I case. Then under the condition that $(\beta_{0}(wa_{-1}w^{-1})>\tau$ we have
$\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(wa_{-1}w^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{A_{w},a_{w}\}\cdot\dim V_{\beta_{0}}$
$\leq$ $\dim\mathbf{G}-\left(\frac{(b_{1}-b_{2})+(b_{1}-b_{3})+\cdots+(b_{1}-b_{n})}{b_{1}-b_{n}}+\frac{(b_{2}-b_{j})+(b_{3}-b_{j})+\cdots+(b_{j-1}-b_{j})}{b_{1}-b_{n}}\right)$
$\hskip 28.45274pt+\frac{(b_{1}-b_{j})-\tau}{b_{1}-b_{n}}\cdot(n-1)$
$=$ $\dim\mathbf{G}-\frac{(b_{1}-b_{j+1})+(b_{1}-b_{j+2})+\cdots+(b_{1}-b_{n})-(n-j)(b_{1}-b_{j})}{b_{1}-b_{n}}-\frac{\tau}{b_{1}-b_{n}}\cdot(n-1)$
$\leq$ $\dim\mathbf{G}-\frac{(n-1)\tau}{\nu_{0}(a_{1})}.$
Similarly, for $w=(j,2,\dots,j-1,1,j+1,\dots,n-1,n)$ ($12\leq j\leq n-$ 1),
$\mathbf{F}_{w}=\left(\begin{array}[]{ccccccc}1&0&\cdots&0&0&\cdots&*\\ 0&1&\cdots&0&0&\cdots&*\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&1&*&\cdots&*\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&0&0&\cdots&1\end{array}\right)$
and $a_{w}\leq n-1$. So under the condition that $(\beta_{0}(wa_{-1}w^{-1})>\tau$ we have
$\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{A_{w},a_{w}\}\cdot\dim V_{\beta_{0}}$
$\leq$ $\dim\mathbf{G}-\left(\frac{(b_{1}-b_{n})+(b_{2}-b_{n})+\cdots+(b_{n-1}-b_{n})}{b_{1}-b_{n}}+\frac{(b_{j}-b_{j+1})+(b_{j}-b_{j+2})+\cdots+(b_{j}-b_{n-1})}{b_{1}-b_{n}}\right)$
$\hskip 28.45274pt+\frac{(b_{j}-b_{n})-\tau}{b_{1}-b_{n}}\cdot(n-1)$
$=$ $\dim\mathbf{G}-\frac{(b_{1}-b_{n})+(b_{2}-b_{n})+\cdots+(b_{j}-b_{n})-j(b_{j}-b_{n})}{b_{1}-b_{n}}-\frac{\tau}{b_{1}-b_{n}}\cdot(n-1)$
$\leq$ $\dim\mathbf{G}-\frac{(n-1)\tau}{\nu_{0}(a_{1})}.$
Let us check a Weyl element $w$ of type III. If $w=(j,2,\dots i-1,1,i+1,\dots,j-1,n,j+1,\dots n-1,i)$ ($12\leq i<j\leq n-$ 1), then
$\beta_{0}(w\cdot a_{-1}\cdot w^{-1})\leq 0$
and there is nothing to prove. If $w=(i,2,\dots i-1,1,i+1,\dots,j-1,n,j+1,\dots,n-1,j)$, then one can compute that
$\mathbf{F}_{w}=\left(\begin{array}[]{ccccccccc}1&0&\cdots&0&0&\cdots&*&\cdots&0\\ 0&1&\cdots&0&0&\cdots&*&\cdots&0\\ \vdots&\vdots&\cdots&\vdots&\vdots&\cdots&\vdots&\cdots&\vdots\\ 0&0&\cdots&1&*&\cdots&*&\cdots&*\\ 0&0&\cdots&0&1&\cdots&*&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&0&0&\cdots&1&\cdots&0\\ \vdots&\vdots&\cdots&\vdots&\vdots&\cdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&0&0&\cdots&0&\cdots&1\end{array}\right)$
and $a_{w}\leq n-1$. Then under the condition that $\beta_{0}(wa_{-1}w^{-1})>\tau$, we have
$\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(wa_{-1}w^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{A_{w},a_{w}\}\cdot\dim V_{\beta_{0}}$
$\leq$ $\dim\mathbf{G}-\left(\frac{(b_{i}-b_{i+1})+(b_{i}-b_{i+2})+\cdots+(b_{i}-b_{n})}{\nu_{0}(a_{1})}+\frac{(b_{1}-b_{j})+(b_{2}-b_{j})+\cdots+(b_{j-1}-b_{j})}{\nu_{0}(a_{1})}-\frac{b_{i}-b_{j}}{\nu_{0}(a_{1})}\right)$
$\hskip 28.45274pt+\frac{(b_{i}-b_{j})-\tau}{\nu_{0}(a_{1})}\cdot(n-1)$
$=$ $\dim\mathbf{G}-\frac{(n-i)b_{i}+b_{1}+\cdots+b_{i}-b_{j}-\cdots-b_{n}-(j-1)b_{j}-n(b_{i}-b_{j})}{\nu_{0}(a_{1})}-\frac{(n-1)\cdot\tau}{\nu_{0}(a_{1})}$
$\leq$ $\dim\mathbf{G}-\frac{(n-1)\tau}{\nu_{0}(a_{1})}.$
Combining type I, type II, type III cases, we conclude that
$\dim_{H}S_{\tau}^{c}\leq\dim\mathbf{G}-\frac{(n-1)\tau}{\nu_{0}(a_{1})}.$
On the other hand, by Theorem 1.3, we know that
$\dim_{H}S_{\tau}^{c}\geq$ $\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-\tau}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}}$
$=$ $\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})\cdot\frac{\tau}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}$
$=$ $\dim\mathbf{G}-\frac{(n-1)\tau}{\nu_{0}(a_{1})}.$
This completes the proof of Theorem 1.7. ∎
Proof of Theorem 1.8.
It is known that there is a complete classification of complex-linear finite-dimensional irreducible representations of $\mathbf{G}=\operatorname{\mathbf{SL}}_{2}$ described as follows. Let $V_{n}$ be the complex vector space of homogeneous polynomials of degree $n$ in two variables $x$ and $y$ ($\dim V_{n}=n+1$). Define $\rho_{n}:\operatorname{\mathbf{SL}}_{2}\to\operatorname{\mathbf{SL}}_{n+1}(V_{n})$ by
$\rho(g)\cdot f\left(\begin{array}[]{c}x\\ y\end{array}\right):=f\left(g^{-1}\cdot\left(\begin{array}[]{c}x\\ y\end{array}\right)\right)$
for any $g\in\operatorname{\mathbf{SL}}_{2}$ and $f\in V_{n}$. Then $\{\rho_{n}\}_{n\in\mathbb{N}}$ consists of all the complex-linear finite dimensional irreducible representations of $\operatorname{\mathbf{SL}}_{2}$ up to equivalence.
Without loss of generality, we may assume that $\rho=\rho_{n}$ for some $n\in\mathbb{N}$ and write $\{a_{t}\}_{t\in\mathbb{R}}\subset\operatorname{\mathbf{SL}}_{2}(\mathbb{R})$ as
$\left\{a_{t}=\left(\begin{array}[]{cc}e^{\alpha t}&0\\ 0&e^{-\alpha t}\end{array}\right):t\in\mathbb{R}\right\}$
for some $\alpha>0$. Let $\mathbf{P}_{0}$ be the lower triangular subgroup in $\mathbf{G}$. The standard basis of $V_{n}$ is $\{x^{n},x^{n-1}y,\cdots,xy^{n-1},y^{n}\}.$ We choose $\mathbb{C}\cdot x^{n}$ to be the highest weight space $V_{\beta_{0}}$ in $V_{n}$ where $\beta_{0}(a_{t})=-n\alpha\cdot t$. The stabilizer $\mathbf{P}_{\beta_{0}}$ of $V_{\beta_{0}}$ is the subgroup
$\mathbf{P}_{\beta_{0}}=\left(\begin{array}[]{cc}*&0\\ &*\end{array}\right)$
which coincides with the minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$ of $\operatorname{\mathbf{SL}}_{2}$, and
$R_{u}(\overline{\mathbf{P}}_{\beta_{0}})=R_{u}(\overline{\mathbf{P}}_{0})=\left(\begin{array}[]{cc}1&*\\ 0&1\end{array}\right).$
The Bruhat decomposition of $\operatorname{\mathbf{SL}}_{2}$ is the following
$\operatorname{\mathbf{SL}}_{2}(\mathbb{R})=\mathbf{P}_{\beta_{0}}\cdot w_{0}\cdot R_{u}(\overline{\mathbf{P}}_{\beta_{0}})\cup\mathbf{P}_{\beta_{0}}\cdot w_{1}$
where $w_{0}$ is the identity element
$w_{1}=\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right)$
and ${}_{\mathbb{Q}}W=\{w_{0},w_{1}\}$.
Now we consider the upper bound of $\dim_{H}S_{\tau}^{c}$. Let $g\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})$ and write
$g=\left(\begin{array}[]{cc}1&t\\ 0&1\end{array}\right).$
Then one can compute that
$\rho(g)\cdot(x^{n},x^{n-1}y,\dots,y^{n})$
$=$ $(x^{n},x^{n-1}y,\dots,y^{n})\cdot\left(\begin{array}[]{cccccc}{n\choose n}(-t)^{0}&0&\cdots&0&\cdots&0\\ {n\choose n-1}(-t)^{1}&{n-1\choose n-1}(-t)^{0}&\cdots&0&\cdots&0\\ \vdots&\vdots&\cdots&\vdots&\cdots&\vdots\\ {n\choose i}(-t)^{i}&{n-1\choose i}(-t)^{i-1}&\cdots&{i\choose i}(-t)^{0}&\cdots&0\\ \vdots&\vdots&\cdots&\vdots&\cdots&\vdots\\ {n\choose 0}(-t)^{n}&{n-1\choose 0}(-t)^{n-1}&\cdots&{i\choose 0}(-t)^{i}&\cdots&{0\choose 0}(-t)^{0}\end{array}\right).$
By Corollary 6.8 and the formula above, we have $a_{w_{0}}=1/n$ for the morphism $\Psi_{w_{0}}$ defined by
$\rho(R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R}))\cdot\{x^{n}\}.$
Let $g$ be a rational element in $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q})$ with
$g=\left(\begin{array}[]{cc}1&\frac{p}{q}\\ 0&1\end{array}\right)\quad(p\in\mathbb{Z},\;q\in\mathbb{N},\;\gcd(p,q)=1).$
Then $\rho(g)\cdot\mathbb{Z}^{n+1}\cap V_{\beta_{0}}\neq\{0\}$ and there exists a primitive integral vector $(k_{n},k_{n-1},\dots,k_{1},l)\in\mathbb{Z}^{n+1}\setminus\{0\}$ such that
($\left(\begin{array}[]{cccccc}{n\choose n}(-\frac{p}{q})^{0}&0&\cdots&0&\cdots&0\\ {n\choose n-1}(-\frac{p}{q})^{1}&{n-1\choose n-1}(-\frac{p}{q})^{0}&\cdots&0&\cdots&0\\ \vdots&\vdots&\cdots&\vdots&\cdots&\vdots\\ {n\choose i}(-\frac{p}{q})^{n-i}&{n-1\choose i}(-\frac{p}{q})^{n-i-1}&\cdots&{i\choose i}(-\frac{p}{q})^{0}&\cdots&0\\ \vdots&\vdots&\cdots&\vdots&\cdots&\vdots\\ {n\choose 0}(-\frac{p}{q})^{n}&{n-1\choose 0}(-\frac{p}{q})^{n-1}&\cdots&{i\choose 0}(-\frac{p}{q})^{i}&\cdots&{0\choose 0}(-\frac{p}{q})^{0}\end{array}\right)\cdot\left(\begin{array}[]{c}l\\ k_{1}\\ \vdots\\ \vdots\\ \vdots\\ k_{n}\end{array}\right)=\left(\begin{array}[]{c}*\\ 0\\ \vdots\\ 0\\ \vdots\\ 0\end{array}\right).$
One can deduce from the equation above that
$l\cdot{n\choose i}\cdot\left(\frac{p}{q}\right)^{n-i}=k_{n-i}\;(\forall 1\leq i\leq n),\quad|l|=q^{n}$
and $d(g)=q^{n}$. According to equation ($\ast$6) in §6, one can compute that $A_{1}=2/n$. Note that for the Weyl element $w_{1}$, $\beta_{0}(w_{1}a_{-1}w_{1}^{-1})<0$. So by Theorem 1.4, we obtain
$\dim_{H}S_{\tau}^{c}\leq\dim\mathbf{G}-1+\frac{n\alpha-\tau}{2\alpha}\cdot\frac{2}{n}=3-\frac{\tau}{\beta_{0}(a_{-1})}.$
On the other hand, by Theorem 1.3, we have
$\dim_{H}S_{\tau}^{c}\geq\dim\mathbf{G}-\frac{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}-\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}}=3-\frac{\tau}{\beta_{0}(a_{-1})}.$
Therefore, we can conclude that $\dim_{H}S_{\tau}^{c}=3-\tau/\beta_{0}(a_{-1})$. ∎
Let us explain how to deduce [17, Corollary 1.3] (or equivalently [17, Theorem 1.2]) from Theorem 1.7. In [17], we consider a regular one-parameter diagonal subgroup $\{a_{t}\}_{t\in\mathbb{R}}$ acting on the homogeneous space $X_{3}=\operatorname{\mathbf{SL}}_{3}(\mathbb{R})/\operatorname{\mathbf{SL}}_{3}(\mathbb{Z})$. According to [17, Definition 1.1], a point $p=g\cdot\operatorname{\mathbf{SL}}_{3}(\mathbb{Z})\in X_{3}$ is Diophantine of type $\tau$ if and only if there exists a constant $C>0$ such that for any $t>0$
$\eta(a_{t}\cdot p)\geq Ce^{-\tau t}$
where $\eta$ is the injectivity radius function on $\operatorname{\mathbf{SL}}_{3}(\mathbb{R})/\operatorname{\mathbf{SL}}_{3}(\mathbb{Z})$. Now let $p=g\operatorname{\mathbf{SL}}_{3}(\mathbb{Z})\in X_{3}$ which is not Diophantine of type $\tau$. Then by definition, for any $\epsilon>0$, there exists $t_{\epsilon}>0$ such that
$\eta(a_{t_{\epsilon}}\cdot p)<\epsilon e^{-\tau t_{\epsilon}}.$
By [37, Corollary 11.18], for sufficiently small $\epsilon>0$, there exists a unipotent element $u_{\epsilon}\in\operatorname{\mathbf{SL}}_{3}(\mathbb{Z})\setminus\{e\}$ such that $g\cdot u_{\epsilon}\cdot g^{-1}\in\operatorname{Stab}(p)=g\operatorname{\mathbf{SL}}_{3}(\mathbb{Z})g^{-1}$ (the stabilizer of $p$) and
$d_{\operatorname{\mathbf{SL}}_{3}}(a_{t_{\epsilon}}g\cdot u_{\epsilon}\cdot g^{-1}a_{-t_{\epsilon}},e)<\epsilon e^{-\tau t_{\epsilon}}$
where $d_{\operatorname{\mathbf{SL}}_{3}}$ is the metric on $\operatorname{\mathbf{SL}}_{3}(\mathbb{R})$ induced by a norm $\|\cdot\|_{\mathfrak{sl}_{3}}$ on the Lie algebra $\mathfrak{sl}_{3}(\mathbb{R})$ of $\operatorname{\mathbf{SL}}_{3}(\mathbb{R})$. Note that $u_{\epsilon}$ is unipotent
$\log u_{\epsilon}=(u_{\epsilon}-I_{3})-\frac{(u_{\epsilon}-I_{3})^{2}}{2}$
and $2\log u_{\epsilon}\in\mathfrak{sl}_{3}(\mathbb{Z})\setminus\{0\}$. Then one can deduce that there exists a constant $\tilde{C}>0$ depending only on $\operatorname{\mathbf{SL}}_{3}$ such that
$\|\operatorname{Ad}_{\operatorname{\mathbf{SL}}_{3}}(a_{t_{\epsilon}}g)(2\log u_{\epsilon})\|_{\mathfrak{sl}_{3}}<\tilde{C}\epsilon e^{-\tau t_{\epsilon}}$
where $\operatorname{Ad}_{\operatorname{\mathbf{SL}}_{3}}$ is the adjoint representation of $\operatorname{\mathbf{SL}}_{3}$. This implies that $g\in S_{\tau}(\operatorname{Ad}_{\operatorname{\mathbf{SL}}_{3}},\{a_{t}\}_{t\in\mathbb{R}})^{c}$ by Definition 1.1.
Conversely, let $g\in\operatorname{\mathbf{SL}}_{3}(\mathbb{R})$ such that $g\in S_{\tau}(\operatorname{Ad}_{\operatorname{\mathbf{SL}}_{3}},\{a_{t}\}_{t\in\mathbb{R}})^{c}$. Then by Definition 1.1, for any $\epsilon>0$, there exists $t_{\epsilon}>0$ such that
$\delta(\operatorname{Ad}_{\operatorname{\mathbf{SL}}_{3}}(a_{t_{\epsilon}}\cdot g)\mathfrak{sl}_{3}(\mathbb{Z}))<\epsilon\cdot e^{-\tau t_{\epsilon}}.$
By [42, Proposition 3.3], for sufficiently small $\epsilon>0$, there exists a nilpotent element $n_{\epsilon}\in\mathfrak{sl}_{3}(\mathbb{Z})\setminus\{0\}$ such that
$\|\operatorname{Ad}_{\operatorname{\mathbf{SL}}_{3}}(a_{t_{\epsilon}}\cdot g)n_{\epsilon}\|_{\mathfrak{sl}_{3}}<\epsilon\cdot e^{-\tau t_{\epsilon}}$
where $\|\cdot\|_{\mathfrak{sl}_{3}}$ is a norm on the Lie algebra $\mathfrak{sl}_{3}(\mathbb{R})$ of $\operatorname{\mathbf{SL}}_{3}(\mathbb{R})$. Note that
$\exp(2n_{\epsilon})=I_{3}+(2n_{\epsilon})+(2n_{\epsilon})^{2}/2\in\operatorname{\mathbf{SL}}_{3}(\mathbb{Z})\setminus\{e\}.$
Then one can deduce that there exists a constant $\tilde{C}>0$ depending only on $\operatorname{\mathbf{SL}}_{3}$ such that
$d_{\operatorname{\mathbf{SL}}_{3}}(a_{t_{\epsilon}}g\cdot\exp(2n_{\epsilon})g^{-1}a_{-t_{\epsilon}},e)<\tilde{C}\cdot\epsilon\cdot e^{-\tau t_{\epsilon}}.$
By [17, Definition 1.1], this implies that $g\operatorname{\mathbf{SL}}_{3}(\mathbb{Z})$ is not Diophantine of type $\tau$. Consequently, Theorem 1.7 and Remark 1.10 imply [17, Corollary 1.3] when $\mathbf{G}=\operatorname{\mathbf{SL}}_{3}$.